### v1

# Semiparametric Skew-Elliptical Distributions For High-Dimensional Graphical Models

Abstract
We propose a semiparametric approach called elliptical skew-(S)KEPTIC for efficiently and robustly estimating non-Gaussian graphical models. Relaxing the assumption of semiparametric elliptical distributions to the family of meta skew-elliptical that accommodates a skewness component, we derive a new estimator which is an extension of the SKEPTIC estimator in Liu et al., 2012a , based on semiparametric Gaussian copula graphical models, to the case of skew-elliptical copula graphical models. Theoretically, we demonstrate that the elliptical skew-(S)KEPTIC estimator achieves robust parametric convergence rates in both graph recovery and parameters estimation. We conduct numerical simulations to prove the reliable graph recovery performance of the elliptical skew-(S)KEPTIC estimator. Finally, the new method is applied to the daily log-returns of the stocks of the S&P500 index and shows better interpretability compared to the Gaussian copula graphical models.
Keywords—High-dimensional statistics, undirected graphical models, elliptical distributions, skewness, robust statistics.
1 Introduction
We consider the problem of estimating high-dimensional undirected graphical models. Given a $p$-dimensional random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$, we aim to estimate the undirected graph $G:=(V,E)$, where the vertices set $V:=\{1,\ldots,p\}$ contains the number of nodes corresponding to the $p$ variables, and the edge set $E$ indicates the conditional independence relationships between $X_{1},\ldots,X_{p}$. Let $X_{i}$ and $X_{j}$ be random variables associated with the multivariate distribution $\boldsymbol{X}$ and define $X_{\setminus{\{i,j}\}}$ the set of variables excluding $X_{i}$ and $X_{j}$. If $X_{i}$ and $X_{j}$ are conditionally independent given $X_{\setminus{\{i,j}\}}$, then there exists no edge between the corresponding nodes $i$ and $j$ in the graph G, denoted as $(i,j)\notin E$.
One reliable approach to estimate the undirected graphical models is the semiparametric nonparanormal (NPN) method proposed in Liu et al., (2009), which relaxes the assumption of Gaussian data, considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that: $f(\boldsymbol{X})=(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim N(0,\Sigma^{0})$, where $\Omega=(\Sigma^{0})^{-1}$. When $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, no edge connects $X_{i}$ and $X_{j}$ if $\Omega_{ij}=0$. Uncovering sparsity is subject to the estimation of the unknown correlation matrix. Liu et al., 2012a propose the SKEPTIC (Spearman/Kendall estimates preempt transformations to infer correlation) estimator for the nonparanormal distribution, a regularized rank-based correlation coefficient estimator based on Kendall’s tau and Spearman’s rho statistics. The estimated correlation matrix is then puggled into shrinkage procedures that reveal the sparsity of the graph.
The nonparanormal distribution exploits a Gaussian-based kernel and belongs to the broad class of meta-elliptical distributions described in Han and Liu, (2014), as well as Liu et al., 2012b , where a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\mathcal{E}C(0,\Sigma^{0},\xi)$ is considered. In particular, Liu et al., 2012b show that the graph $G$ is encoded in the latent generalized partial correlation matrix $\Xi:=-\left[\text{diag}(\Omega)\right]^{-1/2}\Omega\left[\text{diag}(\Omega)\right]^{-1/2}$, thus no edge connects $X_{i}$ and $X_{j}$ if $\Xi_{ij}=0$, which implies $\Omega_{ij}=0$. The main drawback of the meta-elliptical class of distributions is that, even though it embeds skewed distributions, the skewness term does not contribute to reveal the conditional independence structure. For this reason, we introduce the meta skew-elliptical family of distributions defined as $\boldsymbol{X}\sim MS\mathcal{E}(0,\Sigma^{0},\Lambda;f)$ considering a set of functions $f$ such that $f(\boldsymbol{X})\sim SU\mathcal{E}(0,\Sigma^{0},\Lambda,h^{(p+q)},\tau,\Gamma)$ described in Arellano-Valle and Genton, 2010b .
In this paper, we propose a rank-based correlation matrix estimator, which we refer to as elliptical skew-(S)KEPTIC estimator, extending the SKEPTIC estimator to the $MS\mathcal{E}$ family of distributions, adding respect to nonparanormal models the possibility to rely on graphical models that take into account other distributional features such as directional skewness and heavy tails. The main idea is to consider the $\operatorname{SU\mathcal{E}}$ as a combination of the closed skew-Normal distribution introduced in Domìnguez-Molina et al., (2003), which belongs to the class of $\operatorname{SU\mathcal{E}}$ family of distributions, and possesses the stochastic representation defined through Normal distribution and a half-Normal distribution, for which the SKEPTIC estimator of the correlation matrix is defined. The estimated correlation matrix is plugged into existing shrinkage procedures. For instance, Meinshausen and Bühlmann, (2006) study a neighborhood selection approach with the Lasso operator, Banerjee et al., (2008) and Friedman et al., (2008) propose an approach based on penalized likelihood, and Yuan, (2010) and Cai et al., (2011) introduce, respectively, Graphical Dantzig selector and CLIME which are two methods based on linear programming that provide an improvement in terms of theoretical properties respect to the other shrinkage procedures.
We generalize the results in Liu et al., 2012a extending the SKEPTIC estimator to the elliptical skew-(S)KEPTIC estimator by introducing a directional skewness term, encoded in the matrix $\Lambda$, which assigns greater importance to the asymmetry present in data distributions. The proposed estimator is a version of the SKEPTIC estimator conforming to the ${MS}\mathcal{E}$ distribution and embeds two new subclasses of estimators based on Kendall’s tau and Sparman’s rho statistics. While the nonparanormal distribution is a flexible approach to model non-Normal data, the meta skew-elliptical distribution allows us to quantify the asymmetry (and the kurtosis) presented in the data. When the $\operatorname{SU\mathcal{E}}$ distribution reduces to the $\operatorname{CSN}$ distribution, we consider both the statistics for estimating the unknown correlation matrix and we define the elliptical skew-SKEPTIC estimator. Otherwise, we only consider Kendall’s tau statistic, and we define elliptical skew-KEPTIC estimator since Spearman’s rho statistic is not invariant for the class of elliptical distributions (Hult and Lindskog, , 2002).
The proposed framework extends to a greater variety of semiparametric graphical models, introducing the possibility of modeling the combination of directional skewness and heavy tails data and adding flexibility to the nonparanormal-type of data transformations.
We investigate the theoretical properties of the elliptical skew-(S)KEPTIC estimator, and find bounds equal to twice the bounds of the SKEPTIC estimator. The increase in model flexibility justifies the slight reduction in the robustness of the proposed estimators.
We provide numerical studies to support our theory. Through a backward approach, we define the true sparsity pattern given by $\Omega$ and provide the graph recovery using the elliptical skew-(S)KEPTIC estimator which is compared to the SKEPTIC estimator. The elliptical skew-(S)KEPTIC estimator proves an optimal graph recovery, giving statistical efficiency to the proposed method.
Prior studies by Zareifard et al., (2016), Nghiem et al., (2022), and Sheng et al., (2023) have explored graphical models while incorporating skewness into the sparsity pattern estimation. Zareifard et al., (2016) exploit a multivariate $\operatorname{CSN}$ distribution to define the skew Gaussian graphical models which are estimated through a Bayesian approach. Nghiem et al., (2022) propose a novel nodewise regression approach to estimate the graph on data generated from a generalized multivariate skew-Normal distribution. Instead, Sheng et al., (2023) include the shape parameter in the determination of the precision matrix using the skew-Normal distribution of Azzalini, (1985). They build an algorithm that penalizes the likelihood of this distribution in order to estimate $\Omega$ and thus the graph. However, these studies do not emphasize the direct estimation of the correlation matrix while accounting for skewness.
The paper is organized as follows: in Section 2 we recall the background, in Section 3 we define the ${MS}\mathcal{E}$ distribution, which we use to derive the elliptical skew-(S)KEPTIC estimator and in Section 4 we study the theoretical properties of the proposed estimator. In Section 5, we conduct experiments comparing elliptical skew-(S)KEPTIC estimator with the SKEPTIC estimator through numerical comparisons on synthetic data and with an empirical analysis on the S&P 500 stocks data. Section 6 concludes.
2 Background
In this section, we briefly recall the literature preceding our article: the SKEPTIC estimator of Liu et al., 2012a , the semiparametric ellitpical distributions in Liu et al., 2012b and Han and Liu, (2014), in this work, we rely of the meta-elliptical distribution.. The section concludes with the introduction of $\operatorname{SU\mathcal{E}}$ family of distributions in Arellano-Valle and Genton, 2010b , and its link with the undirected graphical models.
2.1 Notation
Here is the notation used throughout the paper. Let $A=\left[A_{ij}\right]\in\mathbb{R}^{p\times p}$ and $\boldsymbol{a}=(a_{1},\ldots,a_{p})\in\mathbb{R}^{p}$. For $1\leq q<\infty$, we define $\|\boldsymbol{a}\|_{q}=\left(\sum_{i=1}^{p}|\boldsymbol{a}|^{q}\right)^{1/q}$, when $q=\infty$: $\|\boldsymbol{a}\|_{\infty}=\max_{1\leq i\leq p}|\boldsymbol{a}_{i}|$. Regarding the $A$ matrix, when $q=1$: $\|A\|_{1}=\max_{1\leq j\leq q}\sum_{i=1}^{p}|A_{ij}|$, and when $q=\infty$, $|A|_{\infty}=\max_{1\leq i\leq q}\sum_{j=1}^{p}|A_{ij}|$. We denote $\|A\|_{\max}=\max_{i,j}|A_{ij}|$. A vector without its $j$-th element is defined: $\boldsymbol{a}_{\backslash j}=(\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{j-1},\boldsymbol{a}_{j+1},\ldots,\boldsymbol{a}_{p})^{\top}\in\mathbb{R}^{p-1}$, and we denote $A_{\backslash i,\backslash j}\in\mathbb{R}^{p-1\times p-1}$ the matrix without the $i$-$th$ row and the $j$-th column. We define: $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ respectively the minimum and the maximum eigenvalue of $A$. $\Omega$ denotes a sparse matrix, and we define the degree of $\Omega$: $\operatorname{deg}(\Omega)=\max_{1\leq i\leq p}\sum_{j=1}^{p}\operatorname{I}(|\Omega_{ij}|\neq 0)$.
2.2 The Nonparanormal SKEPTIC
A random vector $\boldsymbol{X}=(X_{1},...,X_{p})^{\top}$ has a nonparanormal distribution if there exists functions $\{f_{j}\}_{j=1}^{p}$ such that $\boldsymbol{Z}\equiv f(\boldsymbol{X})\sim\operatorname{N}(0,\Sigma^{0})$, where $f(\boldsymbol{X})=(f_{1}(X_{1}),...,f_{p}(X_{p}))^{\top}$, we then write $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$. Further, if the functions $\{f_{j}\}_{j=1}^{p}$ are monotone and differentiable, the joint distribution of $\boldsymbol{X}$ is a Gaussian copula. The focal point is to find a function, $f(\boldsymbol{X})$, that transforms our data into normally distributed data. Suppose we have $n$ observations for the $j$-th variable, $x_{1j},...,x_{nj}$, we compute the rank of each observation, and indicate the rank with $r_{1j},...,r_{nj}$. Let $\Phi(\cdot)$ denote the Gaussian cumulative distribution function, the $f(\boldsymbol{X})$ that transforms the data is:
$\hat{f}_{j}(x_{ij})=\Phi^{-1}(\hat{r}_{ij}),\quad\text{where:}\quad\hat{r}_{ij}=\frac{r_{ij}}{n+1}.$ (1)
We consider the following statistics:
[…]
Both $\hat{\rho}_{ij}$ and $\hat{\tau}_{ij}$ are non-parametric correlations between the empirical realizations of random variables $X_{j}$ and $X_{i}$. Let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, respectively. The cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$ are denoted by $F_{i}$ and $F_{j}$. The population versions of Spearman’s rho and Kendall’s tau are defined as $\rho_{ij}:=\operatorname{Corr}(F_{i}(X_{i}),F_{j}(X_{i}))$ and $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$, respectively. For nonparanormal distributions, there is a connection between Spearman’s rho and Kendall’s tau to the underlying Pearson’s correlation coefficient $\Sigma^{0}_{ij}$. We recall the SKEPTIC estimator.
Lemma 2.1.
(Kruskal, (1958), Liu et al., 2012a ): Assuming $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, we have $\Sigma_{ij}^{0}=2\operatorname{sin}\left(\frac{\pi}{6}\rho_{ij}\right)=\operatorname{sin}\left(\frac{\pi}{2}\tau_{ij}\right)$. From these results, we define two estimators for the unknown correlation matrix $(\Sigma^{0}_{ij})$ that are $\hat{S}^{\rho}=\left[\hat{S}^{\rho}_{ij}\right]$ and $\hat{S}^{\tau}=\left[\hat{S}^{\tau}_{ij}\right]$, where: $\hat{S}^{\rho}_{ij}=2\operatorname{sin}\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ and $\hat{S}^{\tau}_{ij}=\operatorname{sin}\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$,
Spearman’s rho and Kendall’s tau statistics are invariant under strictly increasing transformations of the underlying random variable. This concept will be fundamental in the definition of the elliptical skew-(S)KEPTIC estimator.
2.3 The meta-elliptical distribution
The meta-elliptical distribution introduced in Han and Liu, (2014) relaxes the assumption of elliptically distributed data, in the same way as the nonparanormal of Liu et al., (2009) do for Gaussian distributions. The meta-elliptical family is a strict extension of the nonparanormal distribution.
Definition 2.2.
[…]
Let $\mathcal{R}_{p}:=\{\Sigma\in\mathbb{R}^{p\times p}:\Sigma^{\top}=\Sigma,\text{diag}\left(\Sigma\right)=1,\Sigma\succeq 0\}$, in analogy to the nonparanormal case, the meta-elliptical distribution is defined as follows.
Definition 2.3.
(Han and Liu, (2014)) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has a meta-elliptical distribution, denoted by $\boldsymbol{X}\sim M\mathcal{E}_{p}(\Sigma^{0},\xi;f)$, if there exist a set of univariate monotone functions $f_{1},\ldots,f_{p}$ such that $(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\mathcal{EC}_{p}(0,\Sigma^{0},\xi)$, where $\Sigma^{0}\in\mathcal{R}_{p}$ is defined as the latent generalized correlation matrix.
For the distributions that belong to the meta-elliptical family, Kendall’s tau statistic in Lemma 2.1 can be exploited to estimate the unknown correlation matrix. Nevertheless, the meaning of a missing edge is different with respect to the nonparanormal case. Indeed, when $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$ a missing edge implies conditional independence. On the other hand, when $\boldsymbol{X}\sim M\mathcal{E}(\Sigma^{0},\xi,f)$ a missing edge means conditional linear independence.
The meta-elliptical family contains many useful distributions among which asymmetric distributions. However, the estimate of the correlation matrix does not directly take into account the skewness parameters. We extend the meta-elliptical family giving a direct contribution of the skewness parameters for the conditional independence relations, and thus considering the $\mathcal{E}$ family of distributions.
2.4 $\mathcal{E}$ distributions
We recall the family of $\operatorname{SU\mathcal{E}}$ distributions described in Arellano-Valle and Genton, 2010b . If a random vector has a $\operatorname{SU\mathcal{E}}$ distribution, we write $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,q}(\mu,\Sigma,\Lambda,h^{(p+q)},\tau,\Gamma)$, where $h^{(p+q)}$ is the density generator. We focus on two specifications of the family which are: the unified skew-$t$ ($\operatorname{SUT}$) distribution (Arellano-Valle and Genton, 2010a ), and the unified skew-Normal ($\operatorname{SUN}$) distribution (Arellano-Valle and Azzalini, (2006)). Table 1 shows how the $\operatorname{SU\mathcal{E}}$ distribution reduces into $\operatorname{SUN}$ and $\operatorname{SUT}$ distributions. Formally, we define the $\operatorname{SU\mathcal{E}}$ distribtuion.
Definition 2.4.
(Arellano-Valle and Genton, 2010b ) A continuous p-dimensional random vector $X$ has a multivariate unified skew-elliptical ($\mathcal{E}$) distribution, denoted by $\boldsymbol{X}\sim SU\mathcal{E}_{p,q}(\mu,\Sigma,\Lambda,h^{(p+q)},\tau,\Gamma)$, if its density function at $x\in\mathbb{R}^{p}$ is
$f_{X}(x)=\frac{1}{F_{q}\left(\tau;\Gamma+\Lambda\bar{\Sigma}\Lambda^{T},h^{(q)}\right)}f_{p}\left(y;\mu,\Sigma,h^{(p)}\right)F_{q}\left(\Lambda z+\tau;\Gamma,h_{Q(z)}^{(q)}\right),$ (4)
where $w=\sigma^{-1}(y-\mu),Q(w)=w^{\top}(\Sigma^{0})^{-1}w,\Lambda$ is a $q\times p$ real matrix controlling shape, $\tau\in\mathbb{R}^{q}$ is the extension parameter, $\Gamma$ is a $q\times q$ positive definite correlation matrix, $f_{p}\left(y;\mu,\Sigma,h^{(p)}\right)=$ $|\Sigma|^{-1/2}h^{(p)}(Q(w))$ denotes the density function of an elliptically contoured distribution with location $\mu\in\mathbb{R}^{p}$, positive definite $p\times p$ dispersion matrix $\Sigma$, with $p\times p$ scale and correlation matrices $\sigma=\operatorname{diag}(\Sigma)^{1/2}$ and $\Sigma^{0}=\sigma^{-1}\Sigma\sigma^{-1}$, respectively, and density generator $h^{(p)},F_{r}\left(y;\Sigma,h^{(r)}\right)$ denotes the $r$-dimensional centered elliptical cumulative distribution function with $r\times r$ dispersion matrix $\Sigma$ and density generator $h^{(q)}$, and $h_{Q(w)}^{(q)}(u)=h^{(p+q)}(u+Q(w))/h^{(p)}(Q(w))$.
$\operatorname{SU\mathcal{E}}$ $h^{(p)}$
$\operatorname{SUN}$ $h^{(p)}=\phi^{(p)}$$(u)$where: $\phi^{(p)}(u)=\frac{1}{(2\pi)^{p/2}}\exp\left(-\frac{u^{2}}{2}\right)$
$\operatorname{SUT}$ $h^{(p)}=t^{(p)}(u,\nu)$ where: $t^{(p)}(u,\nu)=\frac{\Gamma\left(\frac{\nu+p}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)\left(\pi a\right)^{p/2}}\left(1+\frac{u}{\nu}\right)^{-\frac{\nu+p}{2}}$
Table 1: Subclasses of skew-elliptical distribution based on the density generator $h^{(p)}$.
The relation between the $\operatorname{SUN}$ and the $\operatorname{SU\mathcal{E}}$ is defined through the following stochastic representation.
Proposition 2.5.
(Arellano-Valle and Genton, 2010b ) Given $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,q}(\mu,\Sigma,\Lambda,h^{(p+q)},\tau,\Gamma)$, when the extension parameter $\tau=0$, we can write $\boldsymbol{X}$ as a mixture of a $\operatorname{SUN}$ distribution: $\boldsymbol{X}=\mu+\sigma V^{-1/2}\boldsymbol{Z}_{0}$, where $V_{0}$ is the cumulative distribution function of a uni-dimensional random variable, called mixture variable, such that: $V_{0}(0)=0$, and $V_{0}$ is independent of $\boldsymbol{Z}_{0}\sim SUN_{p,q}(0,\Sigma^{0},\Lambda,0,\Gamma)$.
The $\operatorname{SUN}$ distribution results in a different parameterization of the closed skew-Normal (CSN) distribution in Domìnguez-Molina et al., (2003). The CSN distribution corresponds to the case in which the cumulative distribution functions in Definition 2.4 are the Gaussian cumulative distribution functions. Therefore, the $\operatorname{SU\mathcal{E}}$ distribution can be expressed in terms of the $\operatorname{CSN}$ distribution. This is useful for deriving the moments and for sampling random numbers from the $\operatorname{SU\mathcal{E}}$ distribution. The $\operatorname{CSN}$ distribution has the following stochastic representation as a combination of a Normal distribution and a half-Normal distribution.
Definition 2.6.
(Domìnguez-Molina et al., (2003)) Let $\boldsymbol{Y}\sim\operatorname{CSN}_{p,q}(\mu,\Sigma,\Gamma,0,\Delta)$, the random vector $\boldsymbol{Y}$ admits the following stochastic representation:
[…]
where: $\boldsymbol{U}\sim\operatorname{N}_{p}(0,I_{p})$ and $\boldsymbol{V}\sim\operatorname{HN}_{q}(0,\Delta+\Gamma\Sigma\Gamma)$, the Normal and the half-Normal distribution, respectively.
The stochastic representation in (5) allows us to estimate an undirected decomposable graph.
Definition 2.7.
(Zareifard et al., (2016)) Consider an undirected decomposable graph $G=(V,E)$, a random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ is called a skew decomposable graphical model concerning the graph G with mean $\mu$, the precision matrix $\Omega$ and the skewness parameter $\alpha\in\mathbb{R}^{p}$ if its density is $\operatorname{CSN}_{p,p}(\mu,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ or equivalently
$f_{\boldsymbol{X}}(x)=2^{p}\phi_{p}\left(x;\mu,\Omega^{-1}\right)\Phi_{p}\left(D_{\alpha}L(x-\mu);0,D^{-1}_{\kappa}\right),$ (6)
where $L$ and $D_{\kappa}$ are the matrix resulting from the modified Cholesky of $\Omega=LD_{\kappa}L^{\top}$, and zero in $L$ matrix determines the conditional independence relations between the $X_{1},\ldots,X_{p}$ variables, further, when the matrix $D_{\alpha}=0$, the random vector $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(\mu,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ reduces to $\boldsymbol{X}\sim\operatorname{N}_{p}(\mu,\Omega^{-1})$.
We introduce an estimator for skew-Gaussian graphical models defined in Zareifard et al., (2016) that extends the SKEPTIC including a skewness component to accommodate non-symmetric data.
3 The elliptical skew-(S)KEPTIC estimator
Definition 3.1.
(Meta skew-elliptical distribution) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has meta skew-elliptical distribution, denoted by $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, if there exists a set of monotone univariate functions $f_{1},\ldots,f_{p}$ and a nonnegative random variable $\xi$, such that
$(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{SU}\mathcal{E}_{p,p}(0,\Sigma^{0},\Lambda,0,\Gamma),$ (7)
where $\Lambda$ is a matrix containing the skewness terms.
Proposition 3.2.
When $\Lambda=0$, then the meta skew-elliptical distribution reduces to the meta-elliptical distribution.
Proposition 3.3.
When $\Lambda=0$ and $h^{(p)}=\phi^{(p)}(u)$, then the meta skew-elliptical distribution reduces to the nonparanornmal distribution.
We exploit the stochastic representation of $\operatorname{SU\mathcal{E}}$ distribution through a CSN distribution (Proposition 2.5) to define the skew-elliptical graphical model. That is, we propose a nonparametric rank-based estimator, which we call elliptical skew-(S)KEPTIC estimator, based on the stochastic representation (5), according to the parametrization of $\operatorname{CSN}$ distribution in (6):
$\boldsymbol{Z}\stackrel{{\scriptstyle p}}{{=}}\boldsymbol{\mu}+L^{-1}D_{\kappa}^{-\frac{1}{2}}D_{\alpha}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{U}+L^{-1}D_{\kappa}^{-\frac{1}{2}}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{V},$ (8)
where: $\boldsymbol{U}\sim N_{p}(0,I_{p})$ and $\boldsymbol{V}\sim HN_{p}(0,I_{p})$, where $N_{p}(0,I_{p})$ is a $p$-dimensional Normal distribution, and $HN_{p}(0,I_{p})$ is a $p$-dimensional half-Normal distribution. The distributions of $\boldsymbol{U}$ and $\boldsymbol{V}$ are independent. The first two moments of $\boldsymbol{Z}$ are:
$\mathbb{E}(\boldsymbol{Z})$ $=\mu+L^{\top}D_{\kappa}^{-\frac{1}{2}}D,$
[…]
where $D=\sqrt{\frac{2}{\pi}}D_{\alpha}(I_{p}+D^{2}_{\alpha})^{-\frac{1}{2}}$.
Lemma 3.4.
[…]
$\boldsymbol{X}\sim\operatorname{SU}\mathcal{E}_{p,p}(\xi,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$ reduces to a $\operatorname{SUT}$ distribution: $\boldsymbol{X}\sim\operatorname{SUT}_{p,p}(\xi,\Sigma,\Lambda,\nu,0,\Gamma)$.
Theorem 3.5 (elliptical skew-KEPTIC estimator).
Assuming $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, such that $\boldsymbol{Z}=f(\boldsymbol{X})$ denotes a generic $\operatorname{SU\mathcal{E}}$ distribution: $\boldsymbol{Z}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma)$. The elliptical skew-KEPTIC estimator is:
${S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
As discussed in Hult and Lindskog, (2002), Spearman’s rho is not invariant in the class of elliptical distributions, and its relation with linear correlation coefficient in Lemma 2.1 holds if and only if the distribution is nonparanonrmal.
Proof.
We consider $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, where $f(\boldsymbol{X})\sim V_{0}^{-1/2}Z_{0}$, and $Z_{0}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ where $D_{\kappa}=I_{p}$, and $V_{0}=1$. We compute $\mathbb{C}orr(X)$ using Kendall’s tau statistic, $\Sigma^{0}_{\tau}=\left[\tau_{ij}\right]$, and we decompose $\Sigma^{0}_{\tau}=L^{\top}D_{\kappa}L$, aligning it $\Sigma^{0}_{\tau}$ to the assumption of $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$ exploiting (9).
[…]
$=(L^{\top}ML)^{-1}\quad\text{where:}\quad M=I-D^{2},$
$=\frac{Adj(L^{\top}ML)}{det(L^{\top}ML)},$
$=\frac{Adj(L^{\top}ML)}{det(L)^{2}det(M)}.$
Where $Adj$ is the adjoint matrix, and $det$ is the determinant of the matrix. The matrices $I$ and $D^{2}$ are both diagonal matrices, and the matrix $M$ is. Thus, we have a new decomposition of the correlation matrix in which the skewness parameter $\alpha$, belonging to the matrix $D_{\alpha}$, contributes to the determination of the linear relations. The matrix $(L^{\top}ML)$ is:
$\left(\begin{array}[]{ccccc}m_{11}&\ell_{12}m_{11}&\ell_{13}m_{11}&\ldots&\ell_{1p}m_{11}\\ \ell_{21}m_{11}&\sum^{p-(p-2)}_{k=1}\ell_{k2}\ell_{2k}m_{kk}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{2k}m_{kk}&\ldots&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{pk}m_{kk}\\ \ell_{31}m_{11}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{3k}m_{kk}&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{3k}m_{kk}&\ldots&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{pk}m_{kk}\\ \vdots&\ldots&\ldots&\ddots&\vdots\\ \ell_{p1}m_{11}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{pk}m_{kk}&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{pk}m_{kk}&\ldots&\sum_{k=1}^{p}\ell_{kp}\ell_{pk}m_{kk}\end{array}\right)$
where the generic $m_{k,k}$ term is:
$m_{kk}=I_{kk}-D^{2}_{kk}=1-\frac{2}{\pi}\frac{\alpha_{kk}^{2}}{1+\alpha_{kk}^{2}}.$
Let $A_{{\tau}}$ denote the final correlation matrix of $X$ with Kendall’s tau statistic. We recall that the stochastic representation (8) of the $\operatorname{CSN}$ is a sum of two comonotic copula functions that have the invariance property, so the SKEPTIC estimator is the same for both the distributions, thus the elliptical skew-KEPTIC estimator is the sum of two SKEPTIC estimators and an estimated matrix, $A_{\tau}$, from $f(\boldsymbol{X})\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$.
When $V_{0}\sim Gamma(\nu/2,\nu/2)$, we have that $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(\xi,\Sigma,\Lambda,\nu,0,\Gamma)$ or equivalently: $f(\boldsymbol{X})\sim\frac{Z_{0}}{\sqrt{V_{0}}}$. Since the positive support of Gamma random variable, and the fact that Kendall’s tau is invariant under monotone transformations, we state that the skew-KEPTIC estimator maintains the same ordinal measures of dependencies of CSN distribution even when $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Sigma^{0},\Lambda,\nu,0,\Gamma)$, where $\Gamma=I_{p}$. ∎
Lemma 3.6.
(skew-Normal SKEPTIC estimator) Assuming $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, such that $\boldsymbol{Z}=f(\boldsymbol{X})$ denotes a generic $\operatorname{SU\mathcal{E}}$ distribution, when $\boldsymbol{Z}\sim\operatorname{SU}\mathcal{E}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma)$ reduces to a $\operatorname{CSN}$ distribution, the skew-Normal SKEPTIC estimator is:
${S}^{\rho,\alpha}_{ij}=\begin{cases}4\sin\left(\frac{\pi}{6}A_{{\rho}_{ij}}\right)&\text{if }i\neq j\\ 1&\text{if }i=j\end{cases},\quad\quad{S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
Proof.
We recall Lemma 2.1, where for the Normal distribution, and for $i\neq j$, we have: ${S}_{ij}=2\sin\left(\frac{\pi}{6}{\rho}_{ij}\right)=\sin\left(\frac{\pi}{2}{\tau}_{ij}\right)$. Let $A_{\rho_{ij}}$ denote the final correlation matrix of $\boldsymbol{X}$ with Spearman’s rho statistic, as in the Normal distribution, we have: ${S}^{\rho,\alpha}_{ij}\equiv 4\sin\left(\frac{\pi}{6}A_{{\rho}_{ij}}\right)=2\sin\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)$. ∎
We generally refer to the estimators in Theorem 3.5 and Lemma 3.6 as elliptical skew-(S)KEPTIC estimators. The specification of the inclusion of the Spearman’s rho will be clear in the application. To obtain the correlation matrix using the elliptical skew-(S)KEPTIC estimator, we need to estimate the parameters of the $\operatorname{CSN}$ distribution. The estimation procedure is based on the EM algorithm of Abe et al., (2021) [^1] The estimated parameters come from the skew-Normal distribution of Azzalini and Capitanio, (1999), with an adequate parametrization explained in Arellano-Valle and Azzalini, (2006) we conform the parameters to a CSN distribution..
3.1 Estimating Marginal Distributions
We find estimators $\{f_{j}\}_{j=1}^{p}$ such that the transformed data are distributed according to a $\operatorname{CSN}$. Let $\Psi$ denote the cumulative distribution function of the $\operatorname{CSN}$ distribution, then the estimators are:
$\hat{f}_{csn,j}(x_{ij})=\Psi^{-1}(\hat{r}_{ij}),\quad\text{where: }\hat{r}_{ij}=\frac{r_{ij}}{n+1}.$ (11)
The functions $\{f_{csn,j}(x_{ij})\}_{j=1}^{p}$ are monotone and differentiable, and Gupta et al., (2004) show that $\operatorname{CSN}$ distribution is closed under conditioning therefore, according to Sklar’s theorem, (6) is a $\operatorname{CSN}$ copula.
3.2 Methods for the Estimation of the Precision Matrix and the Undirected Graph
Upon completing the estimation of the correlation matrix, we follow Liu et al., 2012a and estimate the precision matrix capturing the interrelationships among variables within an undirected graph.
3.2.1 The Dantzig selector
The Dantzig selector estimator (Yuan, , 2010) exploits the relation between the coefficients $\theta$ stemming from a multivariate linear regression and the elements of the precision matrix. Sparse precision matrix implies sparse regression coefficients estimation that relies on an algorithm that depends on the tuning parameter $\delta$. Dantzig procedure involves two steps:
  • Estimation: for $j=1,\ldots,p$, calculate
    $\hat{\theta}^{j}$ $=\text{arg}\min_{\theta\in\mathbb{R}^{p-1}}\|\theta\|_{1}\quad\text{subject to }\quad\|\hat{S}_{\backslash j,j}-\hat{S}_{\backslash j,\backslash j}\theta\|_{\infty}\leq\delta,$ (12)
    $\hat{\Omega}_{jj}$ $=\left[1-2(\hat{\theta}^{j})\hat{S}_{\backslash j,j}+(\hat{\theta}^{j})^{\top}\hat{S}_{\backslash j,\backslash j}\hat{\theta}^{j}\right]\quad\text{and }\quad\hat{\Omega}_{\backslash j.j}=-\hat{\Omega}_{\backslash jj}\hat{\omega}^{j}.$
  • Symmetrization:
    $\hat{\Omega}^{gDS}=\text{arg}\min_{\Omega=\Omega^{T}}\|\Omega-\hat{\Omega}\|_{1}.$ (13)
3.2.2 The Glasso Algorithm
Glasso algorithm is developed by Friedman et al., (2008) and it consists in adding a $L_{1}$-norm penalty term in the maximum likelihood estimation of $\Omega$ as follows:
$\Omega^{glasso}=arg\max_{\Omega\succ 0}\left\{\text{log}\ \text{det}\Omega-\text{tr}(S\Omega)-\lambda_{G}\|\Omega\|_{1}\right\}.$ (14)
The penalty term, $\lambda_{G}$, is the parameter that controls the sparsity of the estimated matrix, i.e., the number of zeros in the precision matrix. Banerjee et al., (2008) shows that the shrinking procedure based on the tuning parameter $\lambda_{G}$ is directly proportional to the input covariance matrix. This means that greater values of $\lambda_{G}$ imply a greater presence of zeros in the estimated $\Omega$, therefore a minor number of interconnections are unveiled.
3.2.3 The CLIME Algorithm
CLIME stands for constrained $\ell_{1}$-minimization for inverse matrix estimation. This algorithm, proposed by Cai et al., (2011), is very attractive for high-dimensional data since the estimator is obtained by combining vector solutions solving a linear program. Let $\Omega$ denote a $p\times p$ precision matrix and $\{\hat{\Omega}_{1}\}$ be the solution set of the optimization problem:
$\text{min}\|\Omega\|_{1}\quad\text{subject to:}$: (15)
$\|\Sigma_{n}\Omega-I\|_{\infty}\leq\lambda_{C},$
where $\lambda_{C}$ is the tuning parameter. The final CLIME estimator of $\Omega_{0}$ is defined by the symmetrization of $\hat{\Omega}_{1}$:
$\hat{\Omega}=(\hat{\omega}_{ij}),\quad\text{where:}$: (16)
$\hat{\omega}_{ij}=\hat{\omega}_{ji}=\hat{\omega}^{1}_{ij}I\{|\hat{\omega}^{1}_{ij}\leq|\hat{\omega}^{1}_{ij}|\}+\hat{\omega}^{1}_{ji}I\{|\hat{\omega}^{1}_{ji}\leq|\hat{\omega}^{1}_{ji}|\}.$
The CLIME estimator has the property that the optimization problem defined in (15) can be decomposed into $p$ optimization problems defined as:
$\min\|\boldsymbol{\beta}\|_{1}\quad\text{subject to}\quad\|\Sigma\boldsymbol{\beta}-\boldsymbol{e}_{i}\|_{\infty}\leq\lambda_{n},$ (17)
where $\boldsymbol{e}_{i}$ is defined as a standard unit vector in $\mathbb{R}^{p}$ with 1 in the $i$-th coordinate and 0 in the all other coordinates and $\boldsymbol{\beta}$ is a vector in $\mathbb{R}^{p}$. Solving (17) $p$ times, we obtain the CLIME estimator: $\{\hat{\Omega}_{1}\}=\{\hat{B}\}:=\{(\hat{\boldsymbol{\beta}}_{1},...,\hat{\boldsymbol{\beta}}_{p})\}$.
4 Theoretical properties
We analyze the theoretical properties of the elliptical skew-(S)KEPTIC estimator. The results suggest that our estimator is reliable when dealing with highly skewed data. As Liu et al., 2012a , our results show an exponential concentration rate in $\|\cdot\|_{\max}$. Moreover, we prove optimal bounds for shrinkage procedures. We show the results for the Dantzig selector, however, they can be easily extended to the CLIME and the Glasso procedures.
Theorem 4.1.
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim\operatorname{MS\mathcal{E}}_{p}(\Sigma^{0},\xi,\Lambda;f)$, restricted to the case $h^{(p)}=\phi^{(p)}(u)$, and let $\hat{S}^{\rho,\alpha}_{ij}=4\sin(\frac{\pi}{6}A_{\hat{\rho,}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-SKEPTIC, then with probability at least $1-1/p^{2}$ we have:
$\|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq 16\pi\sqrt{\frac{logp}{n}}.$ (18)
Theorem 4.2.
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim\operatorname{MS\mathcal{E}}_{p}(\Sigma^{0},\xi,\Lambda;f)$ and let $\hat{S}^{\tau,\alpha}_{ij}=2\sin(\frac{\pi}{2}A_{\hat{\tau}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-(S)KEPTIC estimator, then with probability at least $1-1/p^{2}$ we have:
$\|\hat{S}^{\tau,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq 4\pi\sqrt{\frac{logp}{n}}.$ (19)
Proofs of Theorems 4.1 and 4.2 are in Appendix A. The elliptical skew-(S)KEPTIC estimator achieves convergence in high-dimension, thus it is possible to define the following result.
Theorem 4.3.
When we substitute the estimated matrices $\hat{S}^{\rho,\alpha}$ and $\hat{S}^{\tau,\alpha}$ into the parametric graphical lasso (or the CLIME or graphical Dantzig selector), the estimators provide a slower convergence than the Gaussian case ensuring consistency and graph recovery as a Gaussian model.
Proof.
The proof is based on the observation that the sample correlation matrix $\hat{S}$ is a sufficient statistic for all three shrinkage procedures: the graphical lasso, graphical Dantzig selector, and CLIME. As Yuan, (2010) points out, a sufficient condition to ensure that an estimator is consistent is if there exists some constant $c$, such that $\mathbb{P}\left(\|\hat{S}-\Sigma^{0}\|_{\max}>c\sqrt{\frac{\log p}{n}}\right)\leq 1-\frac{1}{p}$, which can be replaced by (18) and (19) in Theorems 4.1 and 4.2. ∎
The high-dimensional convergence of the elliptical skew-(S)KEPTIC estimator also means that the precision matrix is bounded. Similar results are found by Cai et al., (2011) and Yuan, (2010). In particular, let $\hat{\Omega}^{mse-s}$ denote the inverse correlation matrix using the elliptical skew-(S)KEPTIC estimator and consider the Dantzig selector as the shrinkage procedure to find $\hat{\Omega}$. We introduce a class of inverse matrix defined as $\mathcal{M}_{1}(\kappa,\tau,M):=\{\Omega:\Omega\succ 0,\text{diag}(\Omega^{-1})=1,\|\Omega\|_{1}\leq\kappa,\frac{1}{\tau}\leq\lambda_{\min}(\Omega)\leq\lambda_{\max}(\Omega)\leq\tau,\deg(\Omega)\leq M\}$, where $\kappa,\tau>1$. Given Theorem 4.3, we obtain the following result:
Theorem 4.4.
For $1\leq q<\infty$, there exists a constant $c$ that depends on $\kappa$, $\tau$, $\lambda_{\min}(\Omega)$ and $\lambda_{\max}(\Omega)$, such that
$\sup_{\Omega\in\mathcal{M}_{1}(\kappa,\tau,M)}\|\hat{\Omega}^{mse-s}-\Omega\|_{q}=O_{P}\left(M\sqrt{\frac{\log p}{n}}\right).$ (20)
provided that $\lim_{n\rightarrow\infty}\frac{n}{M^{2}\log d}=\infty$ and $\delta=c\sqrt{\frac{\log d}{n}}$, for sufficiently large $c$, where $\delta$ is the parameter in (12).
Proof.
This result follows Yuan, (2010), which proves that the difference between the estimated precision matrix and the theoretical precision matrix is bounded by $\|\hat{S}-\Sigma^{0}\|_{\max}$, where $\hat{S}$ is the estimated correlation matrix and $\Sigma^{0}$ is the theoretical correlation matrix. In our case, to prove the bounds for the estimated correlation matrix we just replace $\hat{S}$ with $\hat{S}^{\rho,\alpha}$ and $\hat{S}^{\tau,\alpha}$ where the bounds are proved in Theorem 4.1 and 4.2. ∎
5 Experiments
We present the empirical performance of the elliptical skew-(S)KEPTIC methods on both synthetic and real datasets, comparing the proposed method throughout this paper with the SKEPTIC estimator in Liu et al., 2012a . In particular, we consider the following methods for the estimation of graphical models. We generate data from the Normal distribution, the CSN distribution, and the SUT distribution. For each generation, we compute the SKEPTIC and the SKEW-(S)KEPTIC estimator and conduct numerical simulations on synthetic data to show the ability of graph recovery of the proposed estimators. In conclusion, we present an empirical analysis where the methods are applied to the log-returns of the S&P500 index.
5.1 Numerical Simulations
The numerical simulations follow these steps. Using the R package huge (Jiang et al., (2021)), we generate the precision matrix $\Omega=(\Sigma^{0})^{-1}$ using the ”random” structure, in which we randomly generate pairs of conditional dependencies in the off-diagonal elements. We sample $n$ observations from $p$-variate Normal, $\operatorname{CSN}$ and $\operatorname{SUT}$ distribution. We set the location parameter equal to zero for the last two, and $\Sigma^{0}=\Omega^{-1}$. We consider the following cases: $p\in\{100,100,200\}$, and $n\in\{200,100,100\}$. The cases $n=100,p=100$, and $n=100,p=200$ consider the high-dimensional scenarios. Regarding the $\operatorname{CSN}$ parameters, we set $\Delta=1$ and the skewness matrix is generated considering the off-diagonal elements equal to zero, whereas the diagonal matrix is generated from $\operatorname{Uniform}(-0.5,0.5)$. We generate the $\operatorname{SUT}$ distribution with degrees of freedom coming from $\operatorname{Uniform}(2,10)$ [^2] The range is determined by estimating the degrees of freedom of the data on stock returns. . The estimators are then plugged into the ”Glasso” algorithm in (14) to obtain $\Omega$.
To evaluate the accuracy of the methods, we use the receiver operating curve ($\operatorname{ROC}$). We define the false positive rate and true positive rate for a fixed set of tuning parameters $\lambda_{G}$. Let $\hat{G}^{\lambda}_{G}=(V,\hat{E}^{\lambda_{G}})$ be an estimated graph using the proposed procedure and a given value of $\lambda_{G}$. We define the number of false positives for a given $\lambda_{G}$, $\operatorname{FP}(\lambda_{G})$, as the number of edges in $\hat{E}^{\lambda_{G}}$ but not in $E$ and the number of false negatives, $\operatorname{FN}(\lambda_{G})$, as the number of edges in $E$ but not in $\hat{E}^{\lambda_{G}}$. We consider the rate of the number of false negative and false positive:
$\operatorname{FNR}(\lambda_{G}):=\frac{\operatorname{FN}(\lambda_{G})}{|E|},\quad\text{and}\quad\operatorname{FPR}(\lambda_{G}):=\operatorname{FP}(\lambda_{G})/\left[\left(\begin{array}[]{l}d\\ 2\end{array}\right)-|E|\right].$ (21)
To demonstrate the comprehensive performance of the analyzed methods across the entire datasets, we define the ROC curve
$(\operatorname{FNR}(\lambda_{G}),1-\operatorname{FPR}(\lambda_{G})).$ (22)
Graphically, we show averaged ROC curves for all the considered scenarios over 100 trials, plotted using the coordinates in (22). Fig. 1 shows numerical simulations of the average ROC curves of the elliptical skew-(S)KEPTIC estimators and the SKEPTIC estimators, which we employ to estimate $\Omega$ using the Glasso algorithm. The estimators exhibit similar performance, consequently yielding an optimal graph recovery in the simulation results, even when accounting for skewness and kurtosis components in the data, thus giving reliability to the elliptical skew-(S)KEPTIC estimators.
Table 2 presents the quantitative comparisons, showing that regardless of the data generation process or the estimator used, the structure of $\Omega$ can be effectively recovered. Incorporating a skewness component and accounting for heavier tails than those in the NPN distribution do not hinder the recovery of $\Omega$’s structure.
Table 2: Quantitative comparisons of the 4 estimators using different distributions to generate data, and compute the $\Omega$ matrix. The graphs are estimated using the Glasso algorithm. The values are in percentage.
SKEPTIC ($\hat{\rho}$) SKEPTIC ($\hat{\tau}$) SKEW-SKEPTIC ($\hat{\rho}$) SKEW-SKEPTIC ($\hat{\tau}$) $n$ $p$ FPR FNR FPR FNR FPR FNR FPR FNR N DATA 200 100 11.11 74.3 11.2 74.3 11.8 74.7 11.4 74.7 s.e (1.3) (3.5) (1.3) (3.5) (1.4) (3.5) (1.3) (3.5) 100 100 15.5 66.9 15.4 67.0 16.1 67.1 15.7 67.2 s.e (1.6) (2.4) (1.0) (2.4) (1.1) (2.4) (1.1) (2.4) 100 200 13.8 64.0 14.0 64.2 14.9 64.0 14.5 64.1 s.e (1.0) (2.4) (1.0) (2.4) (1.1) (2.4) (1.1) (2.4) CSN DATA 200 100 11.5 73.9 11.5 74.0 12.1 74.1 11.6 73.7 s.e (1.4) (3.5) (1.4) (3.5) (3.5) (3.5) (1.4) (3.5) 100 100 15.6 67.2 15.8 67.3 16.0 65.8 15.3 65.1 s.e (1.6) (3.4) (1.6) (3.4) (1.7) (3.5) (1.6) (3.5) 100 200 13.9 64.0 14.1 64.1 14.3 62.1 13.6 61.6 s.e (1.0) (2.4) (1.0) (2.4) (1.1) (2.5) (1.0) (2.5) SUT DATA 200 100 12.6 72.6 12.8 72.8 - - 14.9 61.0 s.e (1.5) (3.5) (1.5) (3.5) (-) (-) (1.7) (3.3) 100 100 16.8 64.4 17.0 64.5 - - 17.1 64.5 s.e (1.7) (3.3) (1.7) (3.3) (-) (-) (1.7) (3.3) 100 200 14.5 60.7 14.5 61.0 - - 14.9 61.0 s.e (1.0) (2.4) (1.0) (2.4) (-) (-) (1.1) (2.4)
Figure 1: Comparison of average ROC curves, using the Glasso algorithm, of the elliptical skew-(S)KEPTIC estimators with SKEPTIC estimators. Results are determined over 100 trials. We consider, by column, three different data generations. And, by row, three different dataset sizes.
5.2 Empirical Analysis: S&P500 Log-Returns
In the empirical analysis, we study the conditional dependence relations on daily returns of S&P500500500500 companies in the period from $04/01/2016$ to $31/12/2022$ downloaded from Yahoo! Finance, using the proposed elliptical skew-(S)KEPTIC estimator and comparing it with the SKEPTIC. After cleaning the data, the dataset consists of 454 stocks, each counting 1762 observations. The log-returns of the stocks in the S&P500 index are empirically skewed and present heavier tails than the Normal distribution (Cont, (2001)).
Figure 2: Illustration of leptokurtic distributions of daily log-returns, with a relevant skewness component. The stock data are from 04/01/2016 to 31/12/2022.
As a motivating example, we benchmark the analysis of Han and Liu, (2014), reporting the asymmetry of log-returns in Figure 2, where we plot histograms of the returns of three stocks in our sample ”eBay Inc.”, ”Coca-Cola” and ”Wells Fargo” belonging to three different sectors, respectively, Consumer Discretionary, Consumer Staples, and Financials. We appreciate the departure from normality reflected in leptokurtic distribution with a relevant skewness component. Moreover, we investigate the Gaussian assumption of the asset log-returns computing normality tests with different significance levels at 5% and 1%. Table 3 indicates a clear and complete deviation from normality in the stock data. This could be attributed to the presence of exceptional events thus a huge presence of outlier value during the analyzed period, which lead to a departure from the assumption of Normal distribution for log-returns.
Table 3: Normality tests S&P500 log-returns data. This table illustrates the number of stocks that reject the null hypothesis of normality at 1% and 5% levels.
Significance Level Kolmogorov-Smirnov Shapiro-Wilk Lilliefors
1% 454 454 454
5% 454 454 454
To estimate the graph, we determine the tuning parameter $\lambda$ using Extended Bayesian Information Criteria (EBIC) developed in Chen and Chen, (2008):
$\operatorname{EBIC}=-2L+Elog(n)+4\gamma Elog(p).$ (23)
The hyperparameter $\gamma\in\left[0,1\right]$, with greater values of $\gamma$ implying a stronger penalization, $L$ denotes the log-likelihood, $n$ the sample size, $E$ the number of non-zero edges and $p$ the number of nodes. We set $\gamma=1$. The tuning parameter is selected as $\lambda_{G}=0.54$ since it minimizes the scores of the first-order differences in EBIC.
For estimating the undirected graph conforming to the SUT distribution, using the sut Kendall estimator, we need to estimate the degrees of freedom of the log-returns distributions. We follow Arellano-Valle and Genton, 2010a using the maximum likelihood estimation method. In likelihood methods, the underlying assumption is that the observations are independent, therefore we apply the best ARIMA model to the series using the R function auto.arima to filter our returns into a white noise series, and we apply the likelihood method to the filtered series to estimate the degrees of freedom using the command mst.mle from the $R$ package sn developed by Azzalini, (2023).
Figure 3 shows the estimated graphs using npn Kendall, csn Kendall, and sut Kendall estimators.
Figure 3: The resulting graphs using the (a) csn Kendall estimator, (b) sut Kendall estimator, and (c) npn Kendall estimator for S&P 500 stock data from $04/01/2016$ to $31/12/2022$.
The 454 stocks are categorized into 11 Global Industry Classification Standard (GICS) sectors, including Consumer Discretionary (51 stocks), Consumer Staples (33 stocks), Energy (21 stocks), Financials (67 stocks), Health Care (56 stocks), Industrials (69 stocks), Information Technology (59 stocks) Telecommunications Services (17 stocks), Materials (25 stocks), Real Estate (28 stocks), and Utilities (28 stocks). The stocks that belong to the same sector tend to cluster. Nevertheless, csn Kendall and sut Kendall show in better way the grouping of the stocks, thus improving the interpretability of the visualization.
Stocks at the periphery of the network are uncorrelated to the others, during the considered period, thus they are uncorrelated with the movement of the market. Peralta and Zareei, (2016) shows that those stocks have a greater weight in constructing a portfolio using the Markowitz approach. Therefore, in high-dimension, a network approach can be valuable for identifying stocks with the least correlation to the market.
Table 4 shows that introducing the skewness and kurtosis components results in a sparser network estimation compared to the Normal case, thus the elliptical skew-(S)KEPTIC estimator improves the interpretability of the graph.
Table 4: Summary statistics of the estimated networks using Kendall’s tau statistic for the correlation matrix estimation. SK = SKEPTIC, SNK = elliptical skew-SKEPTIC, STK = elliptical skew-KEPTIC. DIM = dimension of the dataset. The notation ”$>$” indicates that one method has a greater number of edges compared to the other.
Network  Edges No. Edges diff.
DIM SK STK SNK SK$>$SNK SK$>$STK SNK$>$STK
S&P 500 454 399 303 252 96 147 51
6 Concluding remarks
We propose a rank-based correlation matrix estimator, named the elliptical skew-(S)KEPTIC estimator which extends to the class of the skew elliptical distribution in estimating the high-dimensional graphical models. The elliptical skew-(S)KEPTIC estimator results for Kendall’s tau in the class of skew-elliptical distribution, and it also encompasses Spearman’s rho when the class of distribution reduces to the skew-Normal. We expand the SKEPTIC estimator in Liu et al., 2012a to the class of semiparametric meta skew-elliptical distribution introduced throughout the paper.
Our analysis is not asymptotic, and the estimator is defined in a way in which we do not explicitly estimate the marginal transformation of the data. Theoretically, the slight reduction in the robustness of the estimator is compensated by the extension of the class of distributions over which we estimate the graphical model.
We illustrate a method for modeling data that conforms to a $\operatorname{MS}\mathcal{E}$ distribution by treating the data generation process as a mixture of Normal and half-Normal distributions. For both the multivariate random variables, the SKEPTIC estimator is defined. Thus, the estimator is a sum of two SKEPTIC estimators and a matrix accounting for skewness. This extension yields a sparser pattern compared to a Gaussian graphical model, resulting in improved interpretability of the resulting graphs. Further, the elliptical skew-(S)KEPTIC estimator shows a slower exponential concentration in $\|\cdot\|_{\max}$ rate than the SKEPTIC estimator. Nevertheless, our findings show a reliable outcome that is useful when dealing with data that exhibit high skewness when it cannot be neglected. Our approach’s effectiveness is validated through optimal graph recoveries in numerical simulations.
Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
A.1 Proof of Theorem 4.1
Proof.
Let $\hat{\rho}_{ij}$ denote the Spearman’s rho estimator of $\Sigma^{0}_{ij}=Corr(X_{i},X_{j})$, the estimator is biased (Zimmerman et al., , 2003).
$\mathbb{E}(\widehat{\rho}_{ij})=\frac{6}{\pi(n+1)}\left[\arcsin\left(\Sigma_{ij}^{0}\right)+(n-2)\arcsin\left(\frac{\Sigma_{ij}^{0}}{2}\right)\right],$ (24)
we rewrite (24) as:
$\Sigma^{0}_{ij}=2\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\text{where: }\quad a_{ij}=\frac{\pi\mathbb{E}(\hat{\rho}_{ij})-2\arcsin(\Sigma^{0}_{ij})}{2(n-2)}.$ (25)
The term $|a_{ij}|\leq\frac{\pi}{n-2}$, therefore for all $n>\frac{6\pi}{t}+2$, we have $|a_{ij}|<\frac{t}{6}$ and thus:
$\mathbb{P}\left(\left|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\right|>t\right)=$ (26)
$=p^{2}\mathbb{P}\left(\left|4\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)-2\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
$=p^{2}\mathbb{P}\left(2\left|2\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)-\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
$=p^{2}\mathbb{P}\left(\left|2\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)-\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\right|>\frac{t}{2}\right)$
$\leq p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})+\frac{6}{\pi}a_{ij}\right|>\frac{3}{\pi}t\right)$
$\leq p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{3}{\pi}t-\left|\frac{6}{\pi}a_{ij}\right|\right)$
$\leq p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{3}{\pi}t-\frac{1}{\pi}t\right)$
$=p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)$
The correlation matrix using Spearman rho, $\hat{\Sigma}_{\rho}$, can be decomposed, using modified Cholesky decomposition as $\hat{\Sigma}_{\rho}=\hat{L}\hat{D}_{\kappa}\hat{L}^{\top}$, where $\hat{D}_{\kappa}=\operatorname{I}_{p}$. In our case, we have: $A_{\hat{\rho}_{ij}}=\hat{L}(I-\hat{D}^{2})\hat{L}^{\top}$. By definition the $k$-th element of $(1-\hat{d}_{kk}^{2})<1$. Thus:
$\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)\leq\mathbb{P}\left(\left|2\left({\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)$ (27)
We prove the bound of the estimator by rewriting Spearman’s rho in terms of the U-statistic and using the Hoeffding inequality. This implies:
$\mathbb{P}\left(\sup_{ij}\left|\hat{S}^{\rho,\alpha}-\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)\leq\exp\left(-\frac{nt^{2}}{16\pi^{2}}\right)$ (28)
A.2 Proof of Theorem 4.2
Proof.
Kendall’s tau is an unbiased estimator: $\mathbb{E}(\hat{\tau}_{ij})=\tau_{ij}$, where $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$ is the population version.
$\mathbb{P}\left(\left|\hat{\tau}_{ij}-\tau_{ij}\right|>t\right)=$
$=\mathbb{P}\left(\left|2\sin\left(\frac{\pi}{2}A_{\hat{\tau}_{ij}}\right)-\sin\left(\frac{\pi}{2}\tau_{ij}\right)\right|>t\right)$
$\leq\mathbb{P}\left(\left|2A_{\hat{\tau}_{ij}}-\tau_{ij}\right|>\frac{2}{\pi}t\right)$
As in the Proof in Theorem 4.1 we have: $(I-\hat{D}^{2})<I$. This implies:
$\mathbb{P}\left(\sup_{ij}|\hat{S}^{\tau,\alpha}-\Sigma^{0}|>t\right)\leq\exp\left(-\frac{nt^{2}}{4\pi^{2}}\right)$ (29)
This result is obtained by exploiting the fact that we can write Kendall’s tau in terms of the U-statistic and using the Hoeffding inequality. ∎