### v2

# Robust Semiparametric Graphical Models with Skew-Elliptical Distributions

Abstract
We propose semiparametric estimators called elliptical skew-(S)KEPTIC for efficiently and robustly estimating non-Gaussian graphical models, relaxing the assumption of semiparametric elliptical distributions to the family of meta skew-elliptical, which accommodates a skewness component. Theoretically, we demonstrate that the elliptical skew-(S)KEPTIC estimators achieve robust convergence rates in both graph recovery and parameters estimation. We conduct numerical simulations that prove the reliable graph recovery performance of the elliptical skew-(S)KEPTIC estimators. Finally, the new method is applied to the daily log-returns of the stocks of the S&P500 index and shows higher sparsity compared to the Gaussian copula graphical models.
Keywords: Elliptical distributions, robust statistics, skewness, and undirected graphical models.
1 Introduction
We consider the problem of estimating undirected graphical models. Given a $p$-dimensional random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$, we aim to estimate the undirected graph $G:=(V,E)$, where the vertices set $V:=\{1,\ldots,p\}$ contains the number of nodes corresponding to the $p$ variables, and the edge set $E$ indicates the conditional independence relationships between $X_{1},\ldots,X_{p}$. Let $X_{i}$ and $X_{j}$ be random variables associated with the multivariate distribution $\boldsymbol{X}$ and define $X_{\setminus{\{i,j}\}}$ the set of variables excluding $X_{i}$ and $X_{j}$. If $X_{i}$ and $X_{j}$ are conditionally independent given $X_{\setminus{\{i,j}\}}$, then there exists no edge between the corresponding nodes $i$ and $j$ in the graph G, denoted as $(i,j)\notin E$.
One reliable approach to estimate the undirected graphical models is the semiparametric nonparanormal (NPN) method proposed in Liu et al., (2009), which relaxes the assumption of Gaussian data, considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that: $f(\boldsymbol{X})=(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{N}(0,\Sigma^{0})$, where $\Omega=(\Sigma^{0})^{-1}$. When $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, no edge connects $X_{i}$ and $X_{j}$ if $\Omega_{ij}=0$. Uncovering sparsity is subject to the estimation of the unknown correlation matrix. Liu et al., 2012a propose the SKEPTIC (Spearman/Kendall estimates preempt transformations to infer correlation) estimator for the nonparanormal distribution, a regularized rank-based correlation coefficient estimator based on Kendall’s tau and Spearman’s rho statistics. The estimated correlation matrix is then plugged into shrinkage procedures that reveal the sparsity of the graph.
The nonparanormal distribution exploits a Gaussian-based kernel and belongs to the broad class of meta-elliptical distributions described in Han and Liu, (2014) and in Liu et al., 2012b , where a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\mathcal{E}C(0,\Sigma^{0},\xi)$ is considered. In particular, Liu et al., 2012b show that the graph $G$ is encoded in the latent partial correlation matrix $\Xi:=-\left[\text{diag}(\Omega)\right]^{-1/2}\Omega\left[\text{diag}(\Omega)\right]^{-1/2}$, thus no edge connects $X_{i}$ and $X_{j}$ if $\Xi_{ij}=0$, which implies $\Omega_{ij}=0$. However, as shown by Rossell and Zwiernik, (2021), the dependence structure of the elliptical distribution differs from that of the Gaussian case. In addition, a major pitfall of the meta-elliptical class of distributions is that, even though it embeds skewed distributions, the skewness term does not contribute to reveal the conditional independence structure.
We propose rank-based correlation matrix estimators, which we refer to as elliptical skew-(S)KEPTIC estimators. These estimators refer to the second moment of the meta skew-elliptical family of distributions, that we introduce in this paper considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\operatorname{SU}\mathcal{E}(0,\Sigma^{0},\Lambda,h^{(p+q)},\tau,\Gamma)$, where the unified skew-elliptical ($\operatorname{SU}\mathcal{E}$) distribution is described in Arellano-Valle and Genton, 2010b . Extending the SKEPTIC estimator to the meta skew-elliptical family of distributions, our methodology allows for a semiparametric estimation of graphical models that take into account distributional features such as directional skewness and heavy tails. The main idea is to consider the unified skew-elliptical distribution as a combination of the closed skew-normal distribution introduced in Domìnguez-Molina et al., (2003), which belongs to the class of $\operatorname{SU\mathcal{E}}$ distributions. The class of unified skew-elliptical distributions, which generalizes the closed skew-normal, enjoys key closure properties under marginalization, conditioning, and linear transformations. These properties enable e.g., to define conjugate priors in probit, tobit, and multinomial probit regression models based on $\operatorname{SU\mathcal{E}}$ distributions, as both the prior and posterior remain within the same family (Durante,, 2019, Karling et al.,, 2024).
The unified skew-elliptical distribution can be written in terms of the closed-skew normal, that possesses a stochastic representation defined through a normal distribution and a half-normal distribution, for which the SKEPTIC estimator of the correlation matrix is defined. The estimated correlation matrix is then plugged into existing shrinkage procedures. For instance, Meinshausen and Bühlmann, (2006) study a neighborhood selection approach with the lasso operator, Banerjee et al., (2008) and Friedman et al., (2008) propose an approach based on penalized likelihood, and Yuan, (2010) and Cai et al., (2011) introduce, respectively, graphical Dantzig selector and CLIME which are two methods based on linear programming that provide an improvement in terms of theoretical properties respect to the other shrinkage procedures.
The proposed estimators generalize the SKEPTIC estimator conforming to the meta skew-elliptical distribution and embed two new subclasses of estimators based on Kendall’s tau and Spearman’s rho statistics. While the nonparanormal distribution is a flexible approach to model non-Gaussian data, the approach based on the meta skew-elliptical distribution allows us to quantify the asymmetry and the kurtosis presented in the data. The elliptical skew-(S)KEPTIC estimators possess skewness terms, encoded in the matrix of skewness parameters $\Lambda$, which assigns greater importance to the asymmetry present in data distributions. When the unified skew-elliptical distribution reduces to the closed skew-normal distribution, we consider both the statistics for estimating the unknown correlation matrix and we define the elliptical skew-SKEPTIC estimator. Otherwise, we only consider Kendall’s tau statistic and we define the elliptical skew-KEPTIC estimator, since Spearman’s rho statistic is not invariant for the class of elliptical distributions (Hult and Lindskog,, 2002). The proposed framework extends to graphical models that incorporate both skewness and heavy tails, adding flexibility to nonparanormal-type data transformations.
We prove that the elliptical skew-(S)KEPTIC estimators have deviation from the true correlation matrix $\Sigma^{0}$ that decreases as the sample size grows, showing a fast exponential rate of convergence in $\|\cdot\|_{\max}$ norm. The numerical analysis supports the theory. Through a backward approach, we define the true sparsity pattern given by $\Omega$ and show the graph recovery using the elliptical skew-(S)KEPTIC estimators, which are compared to the SKEPTIC estimator. The elliptical skew-(S)KEPTIC estimators prove an optimal graph recovery, showing statistical efficiency of the proposed method.
Prior studies by Zareifard et al., (2016), Nghiem et al., (2022), and Sheng et al., (2023) have explored graphical models while incorporating skewness into the sparsity pattern estimation. Zareifard et al., (2016) exploit a multivariate closed skew-normal distribution to define the skew Gaussian graphical models, which are estimated through a Bayesian approach. Nghiem et al., (2022) propose a novel nodewise regression approach to estimate the graph on data generated from a generalized multivariate skew-normal distribution. Instead, Sheng et al., (2023) include the shape parameter in the determination of the precision matrix using the skew-normal distribution of Azzalini, (1985). They develop an algorithm that penalizes the likelihood of this distribution to estimate $\Omega$ and, consequently, the graph. However, these studies do not focus on the direct estimation of the correlation matrix while incorporating skewness.
The paper is organized as follows. In Section 2 we describe the background, including a review of the SKEPTIC estimator and the semiparametric elliptical distribution. In Section 3 we introduce the meta skew-elliptical distribution, which forms the basis to derive the elliptical skew-(S)KEPTIC estimators, whereas in Section 4 we provide some shrinkage procedures to estimate the precision matrix $\Omega$. Theoretical properties of the proposed estimator are discussed in Section 5. Section 6 presents simulation analysis comparing the elliptical skew-(S)KEPTIC estimators with the SKEPTIC estimator on synthetic data, and reports an empirical application to S&P 500 stock returns. Section 7 concludes the paper.
2 Background
2.1 Notation
Here is the notation used throughout the paper. Let $A=\left[A_{ij}\right]\in\mathbb{R}^{p\times p}$ and $\boldsymbol{a}=(a_{1},\ldots,a_{p})\in\mathbb{R}^{p}$. For $1\leq q<\infty$, we define $\|\boldsymbol{a}\|_{q}=\left(\sum_{i=1}^{p}|\boldsymbol{a}|^{q}\right)^{1/q}$, when $q=\infty$: $\|\boldsymbol{a}\|_{\infty}=\max_{1\leq i\leq p}|\boldsymbol{a}_{i}|$. Regarding the $A$ matrix, when $q=1$: $\|A\|_{1}=\max_{1\leq j\leq q}\sum_{i=1}^{p}|A_{ij}|$, and when $q=\infty$, $\|A\|_{\infty}=\max_{1\leq i\leq q}\sum_{j=1}^{p}|A_{ij}|$. We denote $\|A\|_{\max}=\max_{i,j}|A_{ij}|$. A vector without its $j$-th element is defined: $\boldsymbol{a}_{\backslash j}=(\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{j-1},\boldsymbol{a}_{j+1},\ldots,\boldsymbol{a}_{p})^{\top}\in\mathbb{R}^{p-1}$, and we denote $A_{\backslash i,\backslash j}\in\mathbb{R}^{p-1\times p-1}$ the matrix without the $i$-thth row and the $j$-th column. We define: $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ respectively the minimum and the maximum eigenvalue of $A$. $\Omega$ denotes a sparse matrix, and we define the degree of $\Omega$: $\operatorname{deg}(\Omega)=\max_{1\leq i\leq p}\sum_{j=1}^{p}\operatorname{I}(|\Omega_{ij}|\neq 0)$.
2.2 The Nonparanormal SKEPTIC
A random vector $\boldsymbol{X}=(X_{1},...,X_{p})^{\top}$ has a nonparanormal distribution if there exists functions $\{f_{j}\}_{j=1}^{p}$ such that $\boldsymbol{Z}\equiv f(\boldsymbol{X})\sim\operatorname{N}(0,\Sigma^{0})$, where $f(\boldsymbol{X})=(f_{1}(X_{1}),...,f_{p}(X_{p}))^{\top}$, we then write $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$. Further, if the functions $\{f_{j}\}_{j=1}^{p}$ are monotone and differentiable, the joint distribution of $\boldsymbol{X}$ is a Gaussian copula. The focal point is to find a function, $f(\boldsymbol{X})$, that transforms our data into normally distributed data.
Given $n$ observations of the $j$-th variable, $x_{1j},\dots,x_{nj}$, we define the empirical cumulative distribution function (ECDF) $\hat{F}_{j}(w)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(x_{ij}\leq w),$ where $\mathbb{I}(\cdot)$ is the indicator function. The transformation function $\hat{f}_{j}(x_{ij})$ is then defined as $\hat{f}_{j}(x_{ij})=\Phi^{-1}(\hat{F}_{j}(x_{ij}))$, where $\Phi^{-1}(\cdot)$ denotes the quantile normal distribution function. In practice, the ECDF at each observation $x_{ij}$ can be approximated using its normalize rank: $\hat{F}_{j}(x_{ij})\approx\hat{r}_{ij}=\frac{r_{ij}}{n+1}$, where $r_{ij}$ is the rank of $x_{ij}$ among $x_{1j},\dots,x_{nj}$. Thus, the transformation results in
$\hat{f}_{j}(x_{ij})=\Phi^{-1}\left(\hat{r}_{ij}\right).$ (1)
This approach ensures that the transformed data approximate a standard normal distribution while preserving the rank structure of the original data. We consider the following statistics:
$\text{(Spearman's rho)}\ \hat{\rho}_{ij}=\frac{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})(r^{k}_{i}-\bar{r}_{i})}{\sqrt{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})^{2}\sum_{k=1}^{n}(r^{k}_{i}-\bar{r}_{i})^{2}}},$ (2)
$\text{(Kendall's tau)}\ \hat{\tau}_{ij}=\frac{2}{n(n-1)}\sum_{1\leq k<k^{\prime}\leq n}\text{sign}((x^{k}_{j}-x^{k^{\prime}}_{j})(x^{k}_{i}-x^{k^{\prime}}_{i})).$ (3)
Both $\hat{\rho}_{ij}$ and $\hat{\tau}_{ij}$ are non-parametric correlations between the empirical realizations of random variables $X_{j}$ and $X_{i}$. Let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, $F_{i}$ and $F_{j}$ denote the cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$, the population versions of Spearman’s rho and Kendall’s tau are defined as: $\rho_{ij}:=\operatorname{Corr}(F_{i}(X_{i}),F_{j}(X_{i}))$ and $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$, respectively. For nonparanormal distributions, there is a connection between Spearman’s rho and Kendall’s tau to the underlying Pearson’s correlation coefficient $\Sigma^{0}_{ij}$. We recall the SKEPTIC estimator.
Lemma 2.1.
(Kruskal, (1958), Liu et al., 2012a ): Assuming $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, we have $\Sigma_{ij}^{0}=2\operatorname{sin}\left(\frac{\pi}{6}\rho_{ij}\right)=\operatorname{sin}\left(\frac{\pi}{2}\tau_{ij}\right)$.
From these results two estimators of the unknown correlation matrix $(\Sigma^{0}_{ij})$ are $\hat{S}^{\rho}=\left[\hat{S}^{\rho}_{ij}\right]$ and $\hat{S}^{\tau}=\left[\hat{S}^{\tau}_{ij}\right]$, where: $\hat{S}^{\rho}_{ij}=2\operatorname{sin}\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ and $\hat{S}^{\tau}_{ij}=\operatorname{sin}\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$. Spearman’s rho and Kendall’s tau statistics are invariant under strictly increasing transformations of the underlying random variable.
2.3 The meta-elliptical distribution
The meta-elliptical distribution introduced in Han and Liu, (2014) relaxes the assumption of elliptically distributed data, in a similar way as the nonparanormal of Liu et al., (2009) does for Gaussian distributions. The meta-elliptical family is a strict extension of the nonparanormal distribution.
Definition 2.2.
(Han and Liu, (2014), Fang, (2018)) A random vector $\boldsymbol{X}$ has an elliptical distribution, $\boldsymbol{X}\sim\mathcal{E}C_{p}(\mu,\Sigma,\xi)$, if $\boldsymbol{X}$ has the following stochastic representation: $\boldsymbol{X}\overset{d}{=}\mu+\xi AU$. Where: $\mu\in\mathbb{R}^{p}$, $q:=rank(A)$, AA is a $p\times q$ matrix such that $\Sigma=AA^{\top}$, $\xi\geq 0$ is a random variable independent of $U$, $U\in\mathbb{S}^{q-1}$, which is uniformly distributed of the unit sphere in $\mathbb{R}^{q}$.
Let $\mathcal{R}_{p}:=\{\Sigma\in\mathbb{R}^{p\times p}:\Sigma^{\top}=\Sigma,\text{diag}\left(\Sigma\right)=1,\Sigma\succeq 0\}$, in analogy to the nonparanormal case, the meta-elliptical distribution is defined as follows.
Definition 2.3.
(Han and Liu, (2014)) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has a meta-elliptical distribution, denoted by $\boldsymbol{X}\sim M\mathcal{E}_{p}(\Sigma^{0},\xi;f)$, if there exist a set of univariate monotone functions $f_{1},\ldots,f_{p}$ such that $(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\mathcal{EC}_{p}(0,\Sigma^{0},\xi)$, where $\Sigma^{0}\in\mathcal{R}_{p}$ is the latent correlation matrix.
For the distributions that belong to the meta-elliptical family, Kendall’s tau statistic in Lemma 2.1 can be exploited to estimate the unknown correlation matrix. Nevertheless, the meaning of a missing edge is different for the nonparanormal case. Indeed, when $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$ a missing edge implies conditional independence. On the other hand, when $\boldsymbol{X}\sim M\mathcal{E}(\Sigma^{0},\xi,f)$ a missing edge means uncorrelation.
[…]
We recall the family of $\operatorname{SU\mathcal{E}}$ distributions described in Arellano-Valle and Genton, 2010b . If a random vector $\boldsymbol{X}$ has a $\operatorname{SU\mathcal{E}}$ distribution, we write $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, where $h^{(2p)}$ is the density generator. We focus on two specifications of the family, which are: the closed skew-normal (CSN) distribution of Domìnguez-Molina et al., (2003) [^1] We build on the closed skew-normal (CSN) distribution, which represents a reparametrization of the unified skew-normal (SUN) distribution employed by Arellano-Valle and Genton, 2010b . and the unified skew-$t$ (SUT) distribution (Arellano-Valle and Genton, 2010a ). Table 1 shows how the $\operatorname{SU\mathcal{E}}$ distribution reduces into CSN and SUT distributions. Formally, we define the $\operatorname{SU\mathcal{E}}$ distribtuion.
Definition 2.4.
(Arellano-Valle and Genton, 2010b, ) A continuous p-dimensional random vector $\boldsymbol{X}$ has a multivariate unified skew-elliptical $\operatorname{(SU\mathcal{E})}$ distribution, denoted by $\boldsymbol{X}\sim SU\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, if its density function at $x\in\mathbb{R}^{p}$ is
$f_{X}(x)=2^{p}f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)F_{p}\left(\Lambda z;\Gamma,h_{Q(z)}^{(p)}\right),$ (4)
where $\mu$ is the location parameter, $\Sigma$ is a $p\times p$ dispersion matrix, $h^{(p)}$ is the density generator, $F_{r}\left(x;\Sigma,h^{(r)}\right)$ the $r$-dimensional centered elliptical cumulative distribution function with $r\times r$ dispersion matrix $\Sigma$ and density generator $h^{(r)}$. Additionally, $z=\sigma^{-1}(x-\mu)$, $\Lambda$ is a $p\times p$ real matrix controlling shape, $Q(z)=z^{\top}(\Sigma^{0})^{-1}z$, $\Gamma$ is a $p\times p$ positive definite symmetric matrix, $f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)=$ $|\Sigma|^{-1/2}h^{(p)}(Q(z))$ denotes the density function of an elliptically contoured distribution with location $\mu\in\mathbb{R}^{p}$. In conclusion, $h_{Q(z)}^{(p)}(u)=h^{(2p)}(u+Q(z))/h^{(p)}(Q(z))$.
$\operatorname{SU\mathcal{E}}$ $h^{(p)}$
CSN $h^{(p)}=\phi^{(p)}$$(u)$  where: $\phi^{(p)}(u)=\frac{1}{(2\pi)^{p/2}}\exp\left(-\frac{u^{2}}{2}\right)$
SUT $h^{(p)}=t^{(p)}(u,\nu)$ where: $t^{(p)}(u,\nu)=\frac{\Gamma\left(\frac{\nu+p}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)\left(\pi a\right)^{p/2}}\left(1+\frac{u}{\nu}\right)^{-\frac{\nu+p}{2}}$
Table 1: Subclasses of skew-elliptical distribution based on the density generator $h^{(p)}$.
The relation between the CSN and the $\operatorname{SU\mathcal{E}}$ is defined through the following stochastic representation.
Proposition 2.5.
(Arellano-Valle and Genton, 2010b ) Given $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, we can write $\boldsymbol{X}$ as a mixture of a CSN distribution: $\boldsymbol{X}=\mu+\sigma V_{0}^{-1/2}\boldsymbol{Z}_{0}$, where $V_{0}$ is the cumulative distribution function of a uni-dimensional random variable, called mixture variable, such that: $V_{0}(0)=0$, and $V_{0}$ is independent of $\boldsymbol{Z}_{0}\sim\operatorname{CSN}_{p,p}(0,\Sigma^{0},\Lambda,0,\Gamma)$.
Lemma 2.6.
(Arellano-Valle and Genton, 2010b, ) If  $\text{ $V_{0}$}\sim Gamma(\nu/2,\nu/2)$, then
$\boldsymbol{X}\sim\operatorname{SU}\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$ reduces to $\operatorname{SUT}$ distribution: $\boldsymbol{X}\sim\operatorname{SUT}_{p,p}(\mu,\Sigma,\Lambda,\nu,0,\Gamma)$.
We construct an undirected graph on the closed skew-normal distribution, following Zareifard et al., (2016).
Definition 2.7.
Consider an undirected decomposable graph $G=(V,E)$, a random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ is called a skew decomposable graphical model concerning the graph G with mean $\mu$, the precision matrix $\Omega$ and the skewness parameter $\alpha\in\mathbb{R}^{p}$ if its density is $\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ or equivalently
$f_{\boldsymbol{X}}(x)=2^{p}\phi_{p}\left(x-\mu;0,\Omega^{-1}\right)\Phi_{p}\left(D_{\alpha}L(x-\mu);0,D^{-1}_{\kappa}\right),$ (5)
where $L$ and $D_{\kappa}$ are the matrix resulting from the modified Cholesky of $\Omega=LD_{\kappa}L^{\top}$, and zero in $L$ matrix determines the conditional independence relations between the $X_{1},\ldots,X_{p}$ variables, in other words: $L_{ij}=0\Leftrightarrow\Omega_{ij}=0$. Further, when the matrix $D_{\alpha}=0$, the random vector $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ reduces to $\boldsymbol{X}\sim\operatorname{N}_{p}(0,\Omega^{-1})$.
The graphical model for the unified skew-elliptical distribution can be constructed by utilizing the sparsity structure of the closed skew-normal distribution. We rely on a semiparametric approach that considers the class of unified skew-elliptical distributions to extend the SKEPTIC estimator of Liu et al., 2012a proposing a robust estimator of graphical models that accounts for skewness and heavy tails.
3 The elliptical skew-(S)KEPTIC estimators
Definition 3.1.
(Meta skew-elliptical distribution) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has meta skew-elliptical distribution, denoted by $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma;f)$, if there exists a set of monotone univariate functions $f_{1},\ldots,f_{p}$ such that
$(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma),$ (6)
where $\Lambda$ is a matrix containing the skewness terms.
Proposition 3.2.
When $\Lambda=0$, then the meta skew-elliptical distribution reduces to the meta-elliptical distribution.
Proposition 3.3.
When $\Lambda=0$ and $h^{(p)}=\phi^{(p)}(u)$, then the meta skew-elliptical distribution reduces to the nonparanornmal distribution.
We employ the stochastic representation of the unified skew-elliptical distribution, characterized by a closed skew-normal distribution $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ (Proposition 2.5), to introduce the elliptical skew-(S)KEPTIC estimators for constructing skew-elliptical graphical models. The closed skew-normal distribution in (5) can be expressed in the following stochastic form:
$\boldsymbol{X}\stackrel{{\scriptstyle d}}{{=}}L^{-1}D_{\kappa}^{-\frac{1}{2}}D_{\alpha}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{U}+L^{-1}D_{\kappa}^{-\frac{1}{2}}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{V},$ (7)
where: $\boldsymbol{U}\sim HN_{p}(0,I_{p})$ and $\boldsymbol{V}\sim N_{p}(0,I_{p})$, where $HN_{p}(0,I_{p})$ is a $p$-dimensional half-normal distribution, and $N_{p}(0,I_{p})$ is a $p$-dimensional normal distribution. The distributions of $\boldsymbol{U}$ and $\boldsymbol{V}$ are independent. The first two moments of $\boldsymbol{X}$ are:
${E}(\boldsymbol{X})$ $=L^{\top}D_{\kappa}^{-\frac{1}{2}}D,$
$\text{cov}(\boldsymbol{X})$ $=\left(L^{\top}D_{\kappa}^{\frac{1}{2}}(I-D^{2})D_{\kappa}^{\frac{1}{2}}L\right)^{-1},$ (8)
where $D=\sqrt{\frac{2}{\pi}}D_{\alpha}(I_{p}+D^{2}_{\alpha})^{-\frac{1}{2}}$.
Theorem 3.4 (elliptical skew-KEPTIC estimator).
Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa})$. The elliptical skew-KEPTIC estimator is:
${S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
As discussed in Hult and Lindskog, (2002), Spearman’s rho is not invariant in the class of elliptical distributions, and its relation with the Pearson correlation coefficient in Lemma 2.1 holds if and only if the distribution is nonparanormal. The proof of the Theorem is in Appendix A.
Lemma 3.5.
(elliptical skew-SKEPTIC estimator) Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution. When $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa})$ reduces to the CSN distribution in (5), the elliptical skew-SKEPTIC estimator is:
${S}^{\rho,\alpha}_{ij}=\begin{cases}4\sin\left(\frac{\pi}{6}A_{{\rho}_{ij}}\right)&\text{if }i\neq j\\ 1&\text{if }i=j\end{cases},\quad\quad{S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
We generally refer to the estimators in Theorem 3.4 and Lemma 3.5 as elliptical skew-(S)KEPTIC estimators. To obtain the correlation matrix using the elliptical skew-(S)KEPTIC estimators, we need to estimate the parameters of the CSN distribution. The estimation procedure is based on the EM algorithm of Abe et al., (2021) [^2] The estimated parameters come from the skew-normal distribution of Azzalini and Capitanio, (1999), with an adequate parametrization explained in Arellano-Valle and Azzalini, (2006) we conform the parameters to a CSN distribution..
4 Methods for the Estimation of the Precision Matrix and the Undirected Graph
4.1 The Dantzig selector
The Dantzig selector estimator (Yuan,, 2010) exploits the relation between the coefficients $\theta$ stemming from a multivariate linear regression and the elements of the precision matrix. A sparse precision matrix implies sparse regression coefficients estimation that relies on an algorithm that depends on the tuning parameter $\delta$. The Dantzig procedure involves two steps:
  • Estimation: for $j=1,\ldots,p$, calculate
[…]
  • Symmetrization:
    $\hat{\Omega}^{gDS}=\text{arg}\min_{\Omega=\Omega^{T}}\|\Omega-\hat{\Omega}\|_{1}.$ (10)
4.2 The Glasso Algorithm
Glasso algorithm is developed by Friedman et al., (2008) and it consists in adding a $L_{1}$-norm penalty term in the maximum likelihood estimation of $\Omega$ as follows:
$\Omega^{glasso}=arg\max_{\Omega\succ 0}\left\{\text{log}\ \text{det}\Omega-\text{tr}(S\Omega)-\lambda_{G}\|\Omega\|_{1}\right\}.$ (11)
The penalty term, $\lambda_{G}$, is the parameter that controls the sparsity of the estimated matrix, i.e., the number of zeros in the precision matrix. Banerjee et al., (2008) shows that the shrinking procedure based on the tuning parameter $\lambda_{G}$ is directly proportional to the input covariance matrix. This means that greater values of $\lambda_{G}$ imply a greater presence of zeros in the estimated $\Omega$, therefore, a minor number of interconnections are unveiled.
4.3 The CLIME Algorithm
CLIME stands for constrained $\ell_{1}$-minimization for inverse matrix estimation. This algorithm, proposed by Cai et al., (2011), is very attractive for high-dimensional data since the estimator is obtained by combining vector solutions by solving a linear program. Let $\Omega$ denote a $p\times p$ precision matrix and $\{\hat{\Omega}_{1}\}$ be the solution set of the optimization problem:
$\text{min}\|\Omega\|_{1}\quad\text{subject to:}$ (12)
$\|\Sigma\Omega-I\|_{\infty}\leq\lambda_{C},$
where $\lambda_{C}$ is the tuning parameter. The final CLIME estimator of $\Omega$ is defined by the symmetrization of $\hat{\Omega}_{1}$:
$\hat{\Omega}=(\hat{\omega}_{ij}),\quad\text{where:}$ (13)
$\hat{\omega}_{ij}=\hat{\omega}_{ji}=\hat{\omega}^{1}_{ij}I\{|\hat{\omega}^{1}_{ij}\leq|\hat{\omega}^{1}_{ij}|\}+\hat{\omega}^{1}_{ji}I\{|\hat{\omega}^{1}_{ji}\leq|\hat{\omega}^{1}_{ji}|\}.$
The CLIME estimator has the property that the optimization problem defined in (12) can be decomposed into $p$ optimization problems defined as:
$\min\|\boldsymbol{\beta}\|_{1}\quad\text{subject to}\quad\|\Sigma\boldsymbol{\beta}-\boldsymbol{e}_{i}\|_{\infty}\leq\lambda_{n},$ (14)
where $\boldsymbol{e}_{i}$ is defined as a standard unit vector in $\mathbb{R}^{p}$ with 1 in the $i$-th coordinate and 0 in the all other coordinates and $\boldsymbol{\beta}$ is a vector in $\mathbb{R}^{p}$. Solving (14) $p$ times, we obtain the CLIME estimator: $\{\hat{\Omega}_{1}\}=\{\hat{B}\}:=\{(\hat{\boldsymbol{\beta}}_{1},...,\hat{\boldsymbol{\beta}}_{p})\}$.
5 Theoretical properties
We investigate the theoretical properties of the elliptical skew-(S)KEPTIC estimators, establishing concentration inequalities in $\|\cdot\|_{\max}$ for our correlation matrix estimators, and the sparsistency of $\hat{\Omega}$ which describes the characteristic where any parameter that is originally zero is correctly estimated as zero, with a probability approaching one. For the rate of convergence of the correlation matrix estimator, we revisit the results of Liu et al., 2012a . For the corresponding precision matrix, we follow the framework introduced in Lam and Fan, (2009) for demonstrating the rate of convergence and the sparsistency of $\hat{\Omega}$. Our findings are based on the application of shrinkage procedures incorporating an $L_{1}$-penalty, with the three methods outlined in Section 4 adhering to this approach.
Theorem 5.1.
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha},h^{(2p)},0,D_{\kappa};f)$, restricted to the case $h^{(p)}=\phi^{(p)}(u)$, and let $\hat{S}^{\rho,\alpha}_{ij}=4\sin(\frac{\pi}{6}A_{\hat{\rho,}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-SKEPTIC, then with probability at least $1-\frac{1}{p}$ we have:
$\|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq\sqrt{\frac{3(2\pi-3)^{2}}{(9-2\pi)^{2}}\frac{\log\left({p}\right)}{n}}.$ (15)
Theorem 5.2.
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha},h^{(2p)},0,D_{\kappa};f)$ and let $\hat{S}^{\tau,\alpha}_{ij}=2\sin(\frac{\pi}{2}A_{\hat{\tau}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-(S)KEPTIC estimators, then with probability at least $1-\frac{1}{p}$ we have:
$\|\hat{S}^{\tau,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq\sqrt{{6\pi^{2}}\frac{\log\left({p}\right)}{n}}.$ (16)
The proofs of Theorems 5.1 and 5.2 are in Appendix A. The elliptical skew-(S)KEPTIC estimators are consistent, so when the sample size $n$ increases, the estimator converges to the true correlation matrix.
Let $\mathcal{S}=\{(i,j):\omega^{0}_{ij}\neq 0\}$ be the support set of off-diagonal nonzero entries, with cardinality $|\mathcal{S}|=s+p$, let $p_{\lambda}(\cdot)$ be a penalty function, and define
$a=\max_{(i,j)\in\mathcal{S}}p^{\prime}_{\lambda}(|\ell_{ij}|),\quad b=\max_{(i,j)\in\mathcal{S}}p^{\prime\prime}_{\lambda}(|\ell_{ij}|).$
where $\ell_{ij}$ is the $(i,j)$ element of the $L$ matrix that stems from $\Omega=L^{\top}ML$. In this work, we consider shrinkage procedures based on $L_{1}$-penalty, which means $a=\lambda$ and $b=0$. We impose the following regularity conditions:
  • (a)
    The eigenvalues of $\Sigma_{0}$ are bounded: $0<\tau_{1}<\lambda_{\min}(\Sigma_{0})\leq\lambda_{\max}(\Sigma_{0})<\tau_{2}<\infty$.
  • (b)
    Assume $a=O\left(\sqrt{\frac{(1+p/(s+1))\log p}{n}}\right)$, $b=o(1)$, and $\min_{(i,j)\in\mathcal{S}}|\omega^{0}_{ij}|/\lambda\to\infty$.
  • (c)
    The penalty is singular at the origin: $\lim_{t\downarrow 0}p_{\lambda}(t)/(\lambda t)=k>0$.
  • (d)
    For some constants $C,D>0$, when $\theta_{1},\theta_{2}<C\lambda$,
    $|p^{\prime\prime}_{\lambda}(\theta_{1})-p^{\prime\prime}_{\lambda}(\theta_{2})|\leq D|\theta_{1}-\theta_{2}|.$
Theorem 5.3.
Suppose that conditions (a), (b), (c), (d) hold. If $\frac{\log p}{n}=O(\lambda^{2})$, then there exists a local minimizer for the maximum likelihood estimator for $\widehat{{M}}$ and an estimate of $\widehat{{L}}$ that satisfy:
$\|\widehat{{M}}-{M}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),\quad\|\widehat{{L}}-{L}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),$
which implies:
$\|{\Omega}-\Omega\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right).$
The proof of Theorem 5.3 is in Appendix A. We know from Lemma 1 in Zareifard et al., (2016) that $L_{ij}=0\Leftrightarrow\Omega_{ij}=0$, so the matrices $L$ and $\Omega$ share the same sparsity patter. For this reason, we provide sparsistency directly on $\Omega$.
Theorem 5.4.
(Sparsistency) Under the conditions (a), (b), (c), (d) for a local minimizer based on $L_{1}$-penalty that satisfies $\|\widehat{\Omega}-\Omega\|_{\max}^{2}=O_{P}\left\{s\log p/n\right\}$ and $\|\widehat{\Omega}-\Omega\|^{2}=O_{P}\left(\eta_{n}\right)$ for a sequence of $\eta_{n}\rightarrow 0$, where $\eta_{n}\equiv s\sqrt{\frac{\log p}{n}}$. If $\log p/n+\eta_{n}=O\left(\lambda_{n1}^{2}\right)$, then with probability tending to $1,\widehat{\omega}_{ij}=0$ for all $(i,j)\in S^{c}$.
The proof of Theorem 5.4 is in Appendix A.
6 Experiments
6.1 Numerical Simulations
We perform numerical simulations by generating synthetic data from the normal distribution, the closed skew-normal distribution, and the unified skew-$t$ distribution. For each dataset, we compute both the SKEPTIC and the skew-(S)KEPTIC estimators, and we evaluate the graph recovery performance of the proposed methods.
The implementation follows these steps. Using the R package huge (Jiang et al., (2021)), we generate the precision matrix $\Omega=(\Sigma^{0})^{-1}$ using the ”random” structure, in which we randomly generate pairs of conditional dependencies in the off-diagonal elements. We sample $n$ observations from $p$-variate normal, CSN and SUT distribution. We set the location parameter equal to zero, and $\Sigma^{0}=\Omega^{-1}$.
We consider the following cases: $p\in\{100,250,300\}$, and $n\in\{200,1000,500\}$. We generate synthetic data assuming $f(\boldsymbol{X})\sim N(0,\Sigma^{0})$, $f(\boldsymbol{X})\sim\operatorname{CSN}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})$, and $f(\boldsymbol{X})\sim\operatorname{SUT}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})\equiv\frac{W}{\sqrt{V_{0}}}$, where $W$ is distributed according to a closed skew-normal distribution, and $V_{0}\sim Gamma(\nu/2,\nu/2)$ as in Proposition 2.5. We set $\nu=3$ to emphasize the heavy tails in the distribution. To generate the data from the closed skew-normal distribution, we rely on the stochastic representation in (7). Moreover, we generate the skewness parameters from $\alpha=0.4\cdot\operatorname{sign}(u-0.5)\cdot\sqrt{|u-1|},\quad\text{where }u\sim\mathcal{U}(0,1)$. These parameters are encoded in the diagonal of the matrix $D_{\alpha}$.
After the data generation, we compute the skew-(S)KEPTIC and the SKEPTIC estimators, which are then plugged into the glasso algorithm in (11) to obtain $\Omega$. The glasso algorithm depends on a tuning parameter $\lambda_{G}$ whose the choice is performed applying the StARS method in Liu et al., (2010), and when the value is found for each dataset we consider a set of 30 tuning parameters in the range $(\lambda_{G}-0.1,\lambda_{G}+0.1)$.
To evaluate the accuracy of the methods, we use the receiver operating curve ($\operatorname{ROC}$). We define the false positive rate and true positive rate for a fixed set of tuning parameters $\lambda_{G}$. When the tuning parameter is determined we estimate the graph: $\hat{G}^{\lambda_{G}}=(V,\hat{E}^{\lambda_{G}})$, to evaluate it, we define the number of false positives for a given $\lambda_{G}$, $\operatorname{FP}(\lambda_{G})$, as the number of edges in $\hat{E}^{\lambda_{G}}$ but not in $E$ and the number of false negatives, $\operatorname{FN}(\lambda_{G})$, as the number of edges in $E$ but not in $\hat{E}^{\lambda_{G}}$. We consider the rate of the number of false negatives and false positives:
$\operatorname{FNR}(\lambda_{G}):=\frac{\operatorname{FN}(\lambda_{G})}{|E|},\quad\text{and}\quad\operatorname{FPR}(\lambda_{G}):=\operatorname{FP}(\lambda_{G})/\left[\left(\begin{array}[]{l}d\\ 2\end{array}\right)-|E|\right].$ (17)
To demonstrate the comprehensive performance of the analyzed methods across the entire datasets, we define the ROC curve
$(\operatorname{FNR}(\lambda_{G}),1-\operatorname{FPR}(\lambda_{G})).$ (18)
Graphically, we show averaged ROC curves for all the considered scenarios over 100 trials, plotted using the coordinates in (18). Fig. 1 shows that the skew-(S)KEPTIC estimators generally outperform the Pearson correlation approach in non-Gaussian and improve upon the standard SKEPTIC estimator. Adjusting for skewness, also for excess kurtosis in the case of the SUT distribution, enhances the estimation of $\Omega$, especially in settings where the distribution presents heavy tails. In some cases, this leads to better recovery compared to the original SKEPTIC estimator, which has been shown to be optimal in Liu et al., 2012a . Table 2 provides a comprehensive overview of the simulation results, presenting quantitative comparisons of various estimators. The average AUC values of the skew-(S)KEPTIC estimators highlight strong performance in graph recovery under non-Gaussian conditions while still maintaining reliability when applied to normally distributed data. These findings suggest that considering estimators that include the skewness and accommodate for heavier tails, extending the nonparanormal distribution, do not hinder the recovery of the $\Omega$ structure. On the contrary, they may offer a practical advantage due to the robustness across different data distributions.
Figure 1: Comparison of average ROC curves, using the glasso algorithm, of the elliptical skew-(S)KEPTIC estimators with SKEPTIC estimators. Results are determined over 50 trials. We consider, by row, three different data generations and, by column, three different dataset sizes.
Table 2: Quantitative comparisons of the 5 correlation matrix estimators using different distributions to generate data, and compute the $\Omega$ matrix. The graphs are estimated using the glasso algorithm. The values are in percentages.
SKEPTIC ($\hat{\rho}$) SKEPTIC ($\hat{\tau}$) SKEW-SKEPTIC ($\hat{\rho}$) SKEW-SKEPTIC ($\hat{\tau}$) PEARSON
$n$ $p$ AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR
N DATA 200 100 85.0 2.06 27.4 85.0 2.07 27.5 92.0 11.7 4.8 92.0 12.8 3.7 86.0 1.8 25.7
s.e (10.0) (2.6) (21.6) (10.0) (2.6) (21.6) (2.0) (7.7) (3.2) (3.0) (7.8) (2.2) (10.0) (2.3) (21.7)
1000 250 98.3 2.7 0.7 98.0 2.7 0.7 93.0 13.2 6.3 90.5 16.7 11.8 98.5 2.4 0.6
s.e (3.0) (5.5) (1.3) (3.0) (5.5) (1.4) (12.5) (24.6) (24.3) (16.0) (31.4) (32.3) (2.4) (4.9) (1.1)
500 300 94.0 1.3 9.8 94.0 1.3 9.7 96.0 7.8 0.3 96.0 8.8 0.1 95.0 1.1 9.4
s.e (7.0) (2.4) (14.4) (7.0) (2.5) (14.2) (4.0) (8.3) (0.5) (4.0) (8.7) (0.2) (7.0) (2.1) (14.1)
CSN DATA 200 100 86.0 5.0 22.8 86.0 5.0 22.2 88.0 17.3 6.3 88.0 18.5 5.7 83.0 4.4 28.9
s.e (10.0) (4.9) (21.1) (10.0) (4.9) (20.8) (4.0) (10.2) (2.1) (5.0) (10.2) (2.1) (10.0) (4.7) (24.1)
1000 250 98.0 1.6 3.2 98.0 1.6 3.0 96.0 8.3 0.7 94.2 11.0 4.5 96.0 1.3 6.7
s.e (3.0) (2.5) (6.6) (3.0) (2.5) (6.2) (5.7) (11.3) (8.1) (10.0) (20.6) (20.7) (6.0) (2.2) (12.5)
500 300 92.0 1.5 14.7 92.1 1.6 14.2 95.0 8.3 1.7 95.0 9.4 0.9 89.5 1.2 19.7
s.e (9.1) (2.2) (19.3) (8.9) (2.2) (18.8) (3.4) (7.8) (1.6) (4.0) (8.3) (0.6) (11.7) (1.8) (24.4)
SUT DATA 200 100 80.0 2.9 37.4 80.0 3.0 36.7 (-) (-) (-) 89.0 15.6 7.1 71.0 2.2 54.9
s.e (10.0) (3.8) (24.2) (10.0) (3.9) (37.4) (-) (-) (-) (2.0) (6.3) (1.7) (12.0) (3.2) (26.8)
1000 250 97.0 2.1 4.7 97.0 2.1 4.5 (-) (-) (-) 96.0 7.7 0.1 89.0 1.7 20.3
s.e (4.0) (4.5) (8.6) (4.0) (4.5) (8.3) (-) (-) (-) (3.7) (7.4) (0.01) (12.0) (4.2) (25.2)
500 300 89.3 1.9 19.8 89.6 1.9 18.9 (-) (-) (-) 94.4 9.7 1.5 80.0 1.6 37.6
s.e (9.1) (3.7) (19.8) (9.0) (3.7) (19.5) (-) (-) (-) (3.3) (6.9) (0.7) (14.4) (3.3) (30.5)
AUC = Area Under Curve, FPR = False Positive Rate, and FNR = False Negative Rate.
6.2 Empirical Analysis: S&P500 Log-Returns
In the empirical analysis, we study the conditional dependence relations on daily returns of S&P500500 companies in the period from $04/01/2016$ to $31/12/2022$ downloaded from Yahoo! Finance, using the proposed elliptical skew-(S)KEPTIC estimators and comparing it with the SKEPTIC. After cleaning the data, the dataset consists of 454 stocks, each counting 1762 observations. The log-returns of the stocks in the S&P500 index are empirically skewed and present heavier tails than the normal distribution (Cont, (2001)).
Figure 2: Illustration of leptokurtic distributions of daily log-returns, with a relevant skewness component. The stock data are from 04/01/2016 to 31/12/2022.
As a motivating example, we benchmark the analysis of Han and Liu, (2014), reporting the asymmetry of log-returns in Figure 2, where we plot histograms of the returns of three stocks in our sample ”eBay Inc.”, ”Coca-Cola” and ”Wells Fargo” belonging to three different sectors, respectively, Consumer Discretionary, Consumer Staples, and Financials. We appreciate the departure from normality reflected in leptokurtic distribution with a relevant skewness component. Moreover, we investigate the Gaussian assumption of the asset log-returns by computing normality tests with different significance levels at 5% and 1%. Table 3 indicates a clear and complete deviation from normality in the stock data. This could be attributed to the presence of exceptional events, thus a huge presence of outlier values during the analyzed period, which lead to a departure from the assumption of normal distribution for log-returns.
Table 3: Normality tests S&P500 log-returns data. This table illustrates the number of stocks that reject the null hypothesis of normality at 1% and 5% levels.
Significance Level Kolmogorov-Smirnov Shapiro-Wilk Lilliefors
1% 454 454 454
5% 454 454 454
To estimate the graphs, given the large number of variables relative to observations we determine the tuning parameter $\lambda_{G}$ using the Extended Bayesian Information Criteria (EBIC) of Chen and Chen, (2008) to ensure a balanced trade-off between model fit and complexity.
$\operatorname{EBIC}=-2L+E\log(n)+4\gamma E\log(p).$ (19)
The hyperparameter $\gamma\in\left[0,1\right]$, with greater values of $\gamma$ implying a stronger penalization, $L$ denotes the log-likelihood, $n$ the sample size, $E$ the number of non-zero edges and $p$ the number of nodes. We set $\gamma=1$, ensuring a strong penalization, as desired. The tuning parameter $\lambda_{G}$ is set equal to $0.54$ the value that minimizes the scores of the first-order differences in EBIC. To align the log-returns to the unified skew-$t$ distribution for computing the elliptical skew-KEPTIC estimator, we estimate the degrees of freedom of the distribution. We follow Arellano-Valle and Genton, 2010a using the maximum likelihood estimation method. In likelihood methods, the underlying assumption is that the observations are independent, therefore, we apply the best ARIMA model to the series using the R function auto.arima to filter our returns into a white noise series, and we apply the likelihood method to the filtered series to estimate the degrees of freedom using the command mst.mle from the R package sn developed by Azzalini, (2023).
Figure 3 shows the estimated graphs using SKEPTIC, elliptical skew-SKEPTIC, and elliptical skew-KEPITC estimators; all the correlation matrix estimators are based on Kendall’s tau statistic.
Figure 3: The resulting graphs of the log-returns of the S&P 500 stocks from $04/01/2016$ to $31/12/2022$. Kendall’s tau statistic is employed to estimate correlation. In the graph: (a) SKEPTIC estimator, (b) elliptical skew-SKEPTIC estimator, and (c) elliptical skew-KEPTIC estimator.
The 454 stocks are categorized into 11 Global Industry Classification Standard (GICS) sectors, including Consumer Discretionary (51 stocks), Consumer Staples (33 stocks), Energy (21 stocks), Financials (67 stocks), Health Care (56 stocks), Industrials (69 stocks), Information Technology (59 stocks) Telecommunications Services (17 stocks), Materials (25 stocks), Real Estate (28 stocks), and Utilities (28 stocks).
The stocks that belong to the same sector tend to cluster, and those at the periphery of the network are uncorrelated with the others during the considered period, thus, they are uncorrelated with the movement of the market. Peralta and Zareei, (2016) show that those stocks have a greater weight in constructing a portfolio using the Markowitz approach. Therefore, in large-dimensions, a network approach can be valuable for identifying stocks with the least correlation to the market.
Table 4 demonstrates that incorporating skewness and kurtosis components leads to a sparser network estimation compared to the normal case.
Table 4: Summary statistics of the estimated networks using Kendall’s tau statistic for the correlation matrix estimation. SK == SKEPTIC, SNK == elliptical skew-SKEPTIC, STK == elliptical skew-KEPTIC. DIM = dimension of the dataset. The notation ”>>” indicates that one method has a greater number of edges compared to the other.
Network  Edges No. Edges diff.
DIM SK STK SNK SK>>SNK SK>>STK SNK>>STK
S&P 500 454 384 284 231 100 153 53
7 Concluding remarks
We propose rank-based correlation matrix estimators, named the elliptical skew-(S)KEPTIC estimators, to provide semiparametric estimation of graphical models built on meta skew-elliptical distributions. The proposed framework includes two classes of estimators: the elliptical skew-KEPTIC estimator which employs Kendall’s tau statistic, and the elliptical skew-SKEPTIC estimator which also uses Spearman’s rho statistic to estimate the correlation matrix.
We study the theoretical properties of the rank-based correlation matrix estimators and establish concentration bounds. These bounds yield statistical rates of convergence under the elementwise maximum norm for both the correlation matrix estimation and the resulting estimation of the precision matrix. We also study the sparsistency of the precision matrix, evaluating the performance of the elliptical skew-(S)KEPTIC estimators in correctly recovering the edges of the graphical model.
The empirical analysis and numerical simulations demonstrate that the elliptical skew-(S)KEPTIC estimators yield the optimal performance when the data deviates from normality. Applied to real data, specifically the log-returns of S&P500 constituent stocks, the resulting graphs are sparser compared to those estimated under normality.
Our findings show a reliable framework that is useful when dealing with data that exhibit high skewness which cannot be neglected, and its effectiveness is validated through the optimal graph recoveries of the numerical simulations.
Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
A.1 Proof of Theorem 3.4
Proof.
We consider $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, where $f(\boldsymbol{X})\sim V_{0}^{-1/2}Z_{0}$, and
$Z_{0}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ where $D_{\kappa}=I_{p}$, and $V_{0}=1$. We compute the correlation matrix of $\boldsymbol{X}$ using Kendall’s tau statistic, $\Sigma^{0}_{\tau}=\left[\tau_{ij}\right]$, and we decompose $\Sigma^{0}_{\tau}=L^{\top}D_{\kappa}L$, aligning it $\Sigma^{0}_{\tau}$ to the assumption of $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$ exploiting (8).
$\text{cov}(\boldsymbol{X})$ $=\left(L^{\top}(I-D^{2})L\right)^{-1},$
$=(L^{\top}ML)^{-1}\quad\text{where:}\quad M=I-D^{2}.$
The matrices $I$ and $D^{2}$ are both diagonal matrices, and the matrix $M$ is. Thus, we have a new decomposition of the correlation matrix in which the skewness parameter $\alpha$, belonging to the matrix $D_{\alpha}$, contributes to the determination of the linear relations. The matrix $(L^{\top}ML)$ is:
$\left(\begin{array}[]{ccccc}m_{11}&\ell_{12}m_{11}&\ell_{13}m_{11}&\ldots&\ell_{1p}m_{11}\\ \ell_{21}m_{11}&\sum^{p-(p-2)}_{k=1}\ell_{k2}\ell_{2k}m_{kk}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{2k}m_{kk}&\ldots&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{pk}m_{kk}\\ \ell_{31}m_{11}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{3k}m_{kk}&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{3k}m_{kk}&\ldots&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{pk}m_{kk}\\ \vdots&\ldots&\ldots&\ddots&\vdots\\ \ell_{p1}m_{11}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{pk}m_{kk}&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{pk}m_{kk}&\ldots&\sum_{k=1}^{p}\ell_{kp}\ell_{pk}m_{kk}\end{array}\right)$
where the generic $m_{k,k}$ term is:
$m_{kk}=I_{kk}-D^{2}_{kk}=1-\frac{2}{\pi}\frac{\alpha_{kk}^{2}}{1+\alpha_{kk}^{2}}.$
Let $A_{{\tau}}\equiv L^{\top}ML$ denote the rescaled final correlation matrix of $\boldsymbol{X}$, estimated using Kendall’s tau statistic. We recall that the stochastic representation (7) of the CSN is a sum of two comonotic copula functions that have the invariance property, so the SKEPTIC estimator is the same for both the distributions, thus the elliptical skew-KEPTIC estimator is the sum of two SKEPTIC estimators and an estimated matrix, $A_{\tau}$, from $f(\boldsymbol{X})\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$.
When $V_{0}\sim Gamma(\nu/2,\nu/2)$, we have that $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Sigma,D_{\alpha}L,\nu,0,D_{\kappa})$ or equivalently: $f(\boldsymbol{X})\sim\frac{Z_{0}}{\sqrt{V_{0}}}$. Since the positive support of Gamma random variable, and the fact that Kendall’s tau is invariant under monotone transformations, we state that the skew-KEPTIC estimator maintains the same ordinal measures of dependencies of the closed skew-normal distribution even when $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Omega^{-1},D_{\alpha}L,\nu,0,D_{\kappa})$. ∎
A.2 Proof of Lemma 3.5
Proof.
We recall Lemma 2.1, where for the normal distribution, and for $i\neq j$, we have: ${S}_{ij}=2\sin\left(\frac{\pi}{6}{\rho}_{ij}\right)=\sin\left(\frac{\pi}{2}{\tau}_{ij}\right)$. Let $A_{\rho}$ denote the final correlation matrix of $\boldsymbol{X}$ with Spearman’s rho statistic, as in the normal distribution, we have: ${S}^{\rho,\alpha}_{ij}\equiv 4\sin\left(\frac{\pi}{6}A_{{\rho}_{t}{ij}}\right)=2\sin\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)$. ∎
A.3 Proof of Theorem 5.1
Proof.
We have: $\hat{S}^{\rho,\alpha}_{ij}=$ $4\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)$ and we introduce $\Sigma^{\rho}$, the correlation matrix using Spearman’s rho statistic. We know that $A_{\rho}=L^{\top}ML$, where $M=I-D^{2}$. By Definition: $\Sigma^{\rho}=L^{\top}D_{\kappa}L$. This implies:
$\left\lVert\widehat{D}_{\kappa}\right\rVert_{\max}\geq\left\lVert\widehat{M}\right\rVert_{\max},$
$\left\lVert L^{\top}\widehat{D}_{\kappa}L\right\rVert_{\max}\geq\left\lVert L^{\top}\widehat{M}L\right\rVert_{\max},$
$\left\lVert\widehat{\Sigma}^{\rho}\right\rVert_{\max}\geq\left\lVert A_{\widehat{\rho}}\right\rVert_{\max},$
$\left\lVert 4\sin\left(\frac{\pi}{6}\widehat{\Sigma}^{\rho}\right)\right\rVert_{\max}\geq\left\lVert 4\sin\left(\frac{\pi}{6}A_{\widehat{\rho}}\right)\right\rVert_{\max}.$
[…]
$E(\widehat{\rho}_{ij})=\frac{6}{\pi(n+1)}\left[\arcsin\left(\Sigma_{ij}^{0}\right)+(n-2)\arcsin\left(\frac{\Sigma_{ij}^{0}}{2}\right)\right].$ (20)
We rewrite (20) as:
$\Sigma^{0}_{ij}=2\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right),\quad\text{where }a_{ij}=\frac{\pi E(\hat{\rho}_{ij})-2\arcsin(\Sigma^{0}_{ij})}{2(n-2)}.$
The term $|a_{ij}|\leq\frac{\pi}{n-2}$. Therefore, for all $n>\frac{6\pi}{t}+2$, we have $|a_{ij}|<\frac{t}{6}$, and thus:
$\text{pr}\left(\left\lVert 4\sin\left(\frac{\pi}{6}\hat{\Sigma}^{\rho}\right)-\Sigma^{0}\right\rVert_{\max}>t\right)$ $=\text{pr}\left(\max_{ij}\left|4\sin\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)-2\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
$=\text{pr}\left(\max_{ij}\left|2\sin\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)-\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right)\right|>\frac{t}{2}\right)$
$\leq\text{pr}\left(\max_{ij}\frac{2\pi}{3}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})-a_{ij}\right|>\frac{t}{2}\right)\quad\text{(Lipshitz function)}$
$\leq\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{3t}{4\pi}-|a_{ij}|\right)$
$\leq\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{3t}{4\pi}-\frac{t}{6}\right)$
$=\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{(9-2\pi)t}{(2\pi-3)}\right)$
$\leq\sum_{ij}\text{pr}\left(\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{(9-2\pi)t}{(2\pi-3)}\right)\quad\text{(Boole's inequality)}$
$\leq p^{2}\exp\left(-\frac{nt^{2}(9-2\pi)^{2}}{(2\pi-3)^{2}}\right)\quad\text{(Hoeffding's inequality)}$
Setting:
$\beta=p^{2}\exp\left(\frac{(9-2\pi)^{2}nt^{2}}{(2\pi-3)^{2}}\right)$
and solving for $t$, we get:
$t=\sqrt{\frac{(2\pi-3)^{2}}{(9-2\pi)^{2}n}\log\left(\frac{p^{2}}{\beta}\right)}$
When $\beta=\frac{1}{p}$, we have $\sqrt{\frac{(2\pi-3)^{2}}{(9-2\pi)^{2}n}\log\left({p^{3}}\right)}$. ∎
A.4 Proof of Theorem 5.2
Proof.
Similarly to the case of Spearman’s rho, we have that $\hat{S}^{\tau,\alpha}_{ij}=$ $2\sin\left(\frac{\pi}{2}A_{\hat{\tau}_{ij}}\right)$. We introduce $\Sigma^{\tau}$, the correlation matrix using Kendall’s tau statistic. We know that $A_{\tau}=L^{\top}ML$, where $M=I-D^{2}$. By Definition: $\Sigma^{\tau}=L^{\top}D_{\kappa}L$. This implies that:
$\left\lVert\widehat{D}_{\kappa}\right\rVert_{\max}\geq\left\lVert\widehat{M}\right\rVert_{\max},$
$\left\lVert L^{\top}\widehat{D}_{\kappa}L\right\rVert_{\max}\geq\left\lVert L^{\top}\widehat{M}L\right\rVert_{\max},$
$\left\lVert\widehat{\Sigma}^{\tau}\right\rVert_{\max}\geq\left\lVert A_{\widehat{\tau}}\right\rVert_{\max},$
$\left\lVert 2\sin\left(\frac{\pi}{2}\widehat{\Sigma}^{\tau}\right)\right\rVert_{\max}\geq\left\lVert 2\sin\left(\frac{\pi}{2}A_{\widehat{\tau}}\right)\right\rVert_{\max}.$
[…]
$E(\hat{\tau}_{ij})=\tau_{ij},$
where
$\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$
is the population version.
$\text{pr}\left(\left\lVert 2\sin\left(\frac{\pi}{2}\hat{\Sigma}^{\tau}\right)-\Sigma^{0}\right\rVert_{\max}>t\right)\quad$ $=\text{pr}\left(\max_{ij}\ \left|2\sin\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)-\sin\left(\frac{\pi}{2}\tau_{ij}\right)\right|>t\right)$
$\leq\text{pr}\left(\max_{ij}\pi\left|\hat{\tau}_{ij}-\tau_{ij}\right|>t\right)\quad\text{(Lipshitz function)}$
$\leq\sum_{ij}\text{pr}\left(\left|\hat{\tau}_{ij}-\tau_{ij}\right|>\frac{t}{\pi}\right)\quad\text{(Boole's inequality)}$
$\leq p^{2}\exp\left(-\frac{nt^{2}}{2\pi^{2}}\right)\quad\text{(Hoeffding's inequality)}$
Setting
$\beta=p^{2}\exp\left(-\frac{nt^{2}}{2\pi^{2}}\right).$
And solving for $t$, we get:
$t=\sqrt{\frac{2\pi^{2}\log\left(\frac{2p^{2}}{\beta}\right)}{n}}.$
When $\beta=\frac{1}{p}$, we have $\sqrt{\frac{2\pi^{2}}{n}\log\left({p^{3}}\right)}.$
A.5 Proof of Theorem 5.3
We start by proving $\|\hat{L}-L\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$ and $\|\hat{M}-M\|_{\max}O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$. Kendall’s tau and Spearman’s rho statistics can be rewritten in terms of U-statistics, and their estimates of correlation matrix achieve $O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$ as rate of convergence. Under assumption (a), we know that $\Sigma^{0}$ has bounded eigenvalues, which implies:
$\|\hat{L}-L\|_{\max}\leq\tau_{2}\|\hat{\Sigma}-\Sigma^{0}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right).$
Regarding the convergence of the estimated $\hat{M}$ matrix, we know that it has diagonal elements defined as: $\hat{m}_{kk}=1-\frac{2}{\pi}\frac{\hat{\alpha}_{kk}}{1+\hat{\alpha}_{kk}}$ where $\hat{\alpha}$ parameters are estimated via the EM algorithm introduced in Abe et al., (2021), for which we assume that the regulation conditions of Wu, (1983) hold. Then $\|\hat{M}-M\|_{\max}O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$. We find bounds for $\|\hat{\Omega}-\Omega\|_{\max}$.
$\|\hat{\Omega}-\Omega\|_{\max}$ $=\|2\sin\left(\frac{\pi}{2}(\hat{L}^{\top}\hat{M}\hat{L}^{\top})\right)^{-1}-2\sin\left(\frac{\pi}{2}({L}^{\top}{M}{L}^{\top})\right)^{-1}\|_{\max}$
$\leq\|(\hat{L}^{\top}\hat{M}\hat{L})^{-1}-({L}^{\top}{M}{L})^{-1}\|_{\max}$
$=\|(\hat{L}\hat{M}^{-1}\hat{L}^{\top})-({L}{M}^{-1}{L}^{\top})\|_{\max}$
$=\|(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}+L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}+LM^{-1}(\hat{L}-L)\|_{\max}$
We define $A_{1}=(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}$, $A_{2}=L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}$, and $A_{3}=LM^{-1}(\hat{L}^{\top}-L^{\top}_{0})$.
$\|A_{1}\|_{\max}=\|(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}\|_{\max}\leq\|\hat{L}-L\|_{\max}\|\hat{M}^{-1}\|_{\max}\|\hat{L}^{\top}\|_{1}$
$\|\hat{L}-L\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),\quad\|\hat{M}^{-1}\|_{\max}=O(1),\quad\|\hat{L}^{\top}\|_{1}=O(s).$
Where $s$ are the non-zero elements in the $L$ matrix.
$A_{2}=\|L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}\|_{\max}\leq\|L\|_{\infty}\|\hat{M}^{-1}-M^{-1}\|_{\max}\|\hat{L}^{\top}\|_{1}$
Where $\|\hat{M}^{-1}-M^{-1}\|_{\max}=\max_{j}\left|\frac{1}{\hat{M}_{jj}}-\frac{1}{M_{0,jj}}\right|=\left|\frac{\hat{M}_{jj}-M_{0,jj}}{\hat{M}_{jj}M_{0jj}}\right|$. This means that: $\|\hat{M}^{-1}-M^{-1}\||_{\max}=O_{P}(\|\hat{M}-M\|_{\max})=O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$.
$\|A_{3}\|_{\max}$ $=\|LM^{-1}(\hat{L}^{\top}-L^{\top}_{0})\|_{\max}$
$\leq\|L\|_{\infty}\|M^{-1}\|_{\infty}\|\hat{L}^{\top}-L^{\top}_{0}\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right)$
We conclude that:
$\|\hat{\Omega}-\Omega\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right).$
A.6 Proof of Theorem 5.4
As discussed in Lam and Fan, (2009), we only need to prove a bound in $\|\hat{\Omega}-\Omega\|_{\max}$ for the three shrinkage procedures proposed to recovery $\Omega$. For the glasso algorithm, the proof follows the proof of Theorem 2, in Lam and Fan, (2009). For the CLIME, the bounds are provided in Theorem 4 in Cai et al., (2011), under the Frobenius norm, which implies $\|\cdot\|_{\max}$. And for the graphical Dantzig selector, the bounds are provided in Section 3 of Yuan, (2010) work.